### v3

# Scaling Multi-Agent Epistemic Planning through GNN-Derived Heuristics

Abstract.
[…]
Key words and phrases:
Multi-Agent Epistemic Planning; Graph-Neural Networks; Learning in Planning; Heuristicsdoi:
QMRY96611. Introduction
Planning scenarios involving multiple interacting entities, referred to as multi-agent, have gained increasing importance due to their relevance in real-world applications, where groups of agents frequently need to interact. However, effectively addressing multi-agent settings poses one of the most interesting challenges in modern AI research: adequately modeling multi-agent interaction while maintaining tractability (BRAFMAN201352). This is because such modeling requires accounting not only for the state of the world, but also for the dynamics of information exchange between agents. Such reasoning, which deals with formalizing belief relationships among multiple agents, is referred to as epistemic reasoning (fagin1994reasoning).
Interest in Multi-agent Epistemic Planning (MEP)—which integrates epistemic reasoning with automated planning—has surged (BBMvD17), and several epistemic planners have been proposed (BolanderA11; muise2015planning; EngesserBMN17; KominisG17; huang2017general; icaps20; KR2021-12; Pham_Son_Pontelli_2023). To the best of our knowledge, only a few systems (icaps20; DBLP:conf/jelia/BuriganaFM23; Pham_Son_Pontelli_2023) are capable of reasoning over this setting without restrictions. Nonetheless, these systems are severely limited by high computational costs, often making solving impractical. This inefficiency stems mainly from two factors: (1) the intrinsic complexity of the underlying representations, which makes applying transitions and evaluating formulas within epistemic states (e-states) substantially harder than in classical planning; and (2) the lack of effective heuristics, which results in a blind, combinatorial search as plan length increases. While the aforementioned works in MEP largely address the first issue, few efforts tackle the latter. A notable exception is the -EFP planner (DBLP:conf/prima/FabianoPSP24), an extension of le2018efp, which integrates heuristics guidance to improve scalability. Our work builds upon this direction, sharing the core objective of designing effective heuristics extraction methods. We argue this focus is essential, as informed search is what enables scalability in planning systems—from classical heuristics planning (BONET20015; 10.5555/1622559.1622565) to Monte Carlo Tree Search (MCTS) in reinforcement learning (RL) (bouzy2006monte).
[…]
The key contributions of this work are as follows:
We define and compare three embeddings for Kripke structures to serve as input to a GNN-based regressor.
We propose a fully automated pipeline for efficient data generation and training of the GNN-based regressor to approximate the perfect heuristics in the MEP setting.
We integrate the GNN-regressor into the MEP solving process, where it is used to evaluate epistemic states by assigning heuristics scores that guide the search process.
We provide a comprehensive evaluation of this integration by thoroughly testing several benchmarks.
These contributions, supported by experimental results, represent a foundational step in integrating ML with MEP.
The remainder of this paper is structured as follows. In Section 2, we provide background on MEP and GNNs. Section 3 presents our main theoretical contribution. In particular, Section 3.1 presents the design of the embedding and the dataset generation while Section 3.4 illustrates the training of the GNN-based regressor. Section 4 reports experimental results that evaluate the performance and scalability of our approach. We discuss limitations and related work in Sections 5 and 6, and conclude in Section 7.
2. Background
2.1. Dynamic Epistemic Logic
[…]
Let us denote as a set of agents such that with , and as a set of propositional variables, referred to as fluents literals, or simply fluents. Each world is described by a subset of elements from intuitively, those deemed True. Furthermore, in epistemic logic, each agent is associated to an epistemic modal operator , signifying the belief222We use the terms knowledge and belief interchangeably, as their distinction is beyond this work’s scope. See fagin1994reasoning for a full discussion. of the agent. Additionally, the epistemic group operator is introduced. Essentially, this operator represents the common knowledge of a group of agents .
To be more precise, as in baral2021action, we have that a fluent formula is a propositional formula built using fluents in as propositional variables and the propositional operators . On the other hand, a belief formula is either
a fluent formula;
if is a belief formula and , then is a belief formula;
if and are belief formulae, then and are belief formulae, where ; or
if is a belief formula and then is a belief formula.
denotes the language of the belief formulae over the set .
The classical way of providing semantics for epistemic logic is in terms of pointed Kripke structures (Kripke1963-KRISCO).
[…]
Let with . A pointed Kripke structure is a pair , such that:
S is a set of worlds;
is a function that associates an interpretation of to each element of S;
for , is a binary relation over S; and
points at the real world.
To elaborate, the component encompasses all the possible worlds configurations, while specifically represents the beliefs held by each individual agent.
Intuitively, to verify whether a belief formula holds, we need to apply reachability within the Kripke model representing the e-state. By exploring the set of reachable worlds obtained by applying epistemic operators, we determine which configurations of fluents an agent (or group of agents) considers possible. Inconsistencies among these reachable worlds are used to model ignorance. The formal semantics over pointed Kripke structures is provided in icaps20; baral2021action and omitted here as it is not integral to understanding the contribution of this paper.
2.2. Multi-Agent Epistemic Planning
We are now ready to introduce the fundamental concepts of MEP, while addressing interested readers to fagin1994reasoning; BolanderA11 for a more exhaustive introduction.
Let us begin by providing the notion of a multi-agent epistemic planning problem in Definition 2.2. Intuitively, an epistemic planning problem encompasses all the necessary information to frame a planning problem within a multi-agent scenario.
[…]
We define a multi-agent epistemic problem as the tuple where:
is the set of all the fluents of ;
is the set of the agents of ;
represents the set of all the actions;
is the set of belief formulae that describes the initial conditions of the planning process; and
is the set of belief formulae that represents the goal conditions.
Note that the tuple captures the domain description of which the problem is an instance.
A solution, or a plan, of a MEP problem is a sequence of actions in that, when executed, transforms the initial e-state into one that satisfies the .
In this context, an epistemic state—represented by a pointed Kripke structure—encapsulates a problem’s “physical” configuration along with the beliefs of the agents.
To the best of our knowledge, the most accepted formalization of a comprehensive action language for multi-agent epistemic planning is (baral2021action). Note that other languages capable of reasoning about DEL also exist (muise2015planning; KR2021-12), but they limit their expressiveness in favor of efficiency and are therefore not considered here.
serves as a high-level action language facilitating reasoning about agents’ beliefs within , where e-states are represented as Kripke structures. It utilizes an English-like syntax, leverages event models to define the transition functions, and uses reachability over Kripke models to characterize entailment.
In their work baral2021action delineate three distinct types of actions within the context of multi-agent domains: (i) world-alteringactions (or ontic actions): employed to modify specific properties, or fluents, within the world—denoted through the statement “ causes 333 can be either a fluent or its negation”; (ii) sensingactions: used by an agent to refine her beliefs about the world—denoted through the statement “ determines ”; and (iii) announcementactions: utilized by an agent to influence the beliefs held by other agents—denoted through the statement “ announces ”. For brevity, we will not provide further details of here. Interested readers can find a comprehensive description in baral2021action.
2.3. Graph Neural Networks (GNNs)
[…]
In our neural estimator, the GNN forms the first stage of the pipeline: it processes the graph representation of an epistemic state and produces a compact latent embedding that serves as the basis for the heuristic estimate, capturing the structural and logical characteristics of the state. Here, message passing is implemented via GINEConv (DBLP:conf/iclr/HuLGZLPL20): given a node and its neighborhood , the update at layer is
[…]
Reviews of architectures and applications are proposed in wu2020comprehensive and zhou2020graph, respectively.
3. Learning MEP Heuristics with GNN
As discussed, planning in MEP is an extremely resource-intensive task. To mitigate this computational burden, we explore alternatives to Breadth-First Search (BFS), such as Best-First Search, which we abbreviate as HFS (for Heuristics-First Search) to distinguish it from BFS. HFS is a widely used strategy that prioritizes expanding states with the most favorable heuristics score. Both BFS and HFS are standard approaches described in modernApproach.
BFS is an uninformed search method, exploring the state space uniformly. In contrast, HFS aims to guide the search more efficiently by prioritizing e-states that are likely closer to the goal. The key challenge lies in defining an effective evaluation function—also known as a heuristics—capable of accurately ranking epistemic states. Heuristics are a central topic in the planning literature and have been extensively studied and proven effective (Keyder:2008:HPA:1567281.1567409; modernApproach). For this reason, a key contribution of this work is the formalization of heuristics tailored for MEP through the employment of GNNs.
3.1. Training Data Generation
The first key challenge in employing ML techniques in MEP is determining what kind of information can be meaningfully extrapolated from data. Several approaches have been proposed in automated planning, such as training models end-to-end using full problem descriptions and their associated goals (ijcai2023p839; huang2024planningdarkllmsymbolicplanning). However, these methods often suffer from low accuracy and require a large number of training instances to generalize effectively (kambhampati2024llmscantplanhelp).
To address these limitations, we adopt a learning approach in which every state represents an independent training signal, rather than relying on complete trajectories. This offers two main advantages. First, although individual predictions may occasionally be inaccurate, their impact on the overall search is limited—as long as the trend of the heuristics is informative, the search remains effective. Second, this approach significantly reduces the amount of data required. Instead of needing complete problems as training instances, we can treat each e-state encountered during exploration as a distinct data point. This allows the generation of tens of thousands of e-states, yielding a large set of training examples in a single run of the planner.
The main objective of this work is to use GNNs to extract information from the data to guide the search. In particular, we aim at approximating the perfect heuristics that assign to each state its distance to the nearest goal.
Illustration of the overall training and inference process.
[…]
3.2. e-State Representation
The next challenge we addressed was designing a suitable data representation for integration within a GNN. Each e-state is formalized as a pointed Kripke structure (Definition 2.1), namely
Encoding the relational structure is straightforward, as each agent can be mapped to a unique integer label. The primary challenge lies in defining a consistent and informative numerical representation of the worlds suitable for GNN input.
We explored three strategies, each offering a different trade-off between information preservation and efficiency. In Section 4.1, we compare these representations, illustrated below, and show that hash-based encoding achieves the best performance.
Symbolic ID Mapping.
Each world is assigned a symbolic integer identifier through a mapping Intuitively, this assigns consecutive integers to newly observed worlds. However, since this assignment depends on the order in which worlds are generated, making not invariant across runs, leading to inconsistencies between datasets. This representation has the lowest computational burden for data generation, training, and inference, but it incurs the greatest information loss, as similarities between worlds are not preserved.
Hash-based Encoding.
A hash function is applied to each world: ; where denotes the repetition index distinguishing identical copies of the same world within an e-state. This ensures consistent world identifiers across runs, but relative distances between world evaluations are lost, as, for example, evaluations differing only by a single fluent value may receive completely unrelated hash values. We use the hash_range function from the Boost libraries (10.5555/2049814) due to its high performance and existing use in deep for e-state storage, which minimizes computational overhead. Alternative hashing approaches may offer further improvements and are left for future work.
This encoding represents a middle ground in terms of the trade-off between information loss and the computational power required.
Bitmask Encoding.
Each world is represented as a purely binary vector: where is the binary representation of the repetition index , and if fluent holds in , and otherwise (or when ). Note that, by construction, is assumed to be greater than . Thus, encodes both the repetition number of the world within the e-state and its truth assignment over as a single binary vector. This embedding preserves most of the logical information about the e-estate, but it comes at the cost of a larger encoding, which requires more data for the model to converge appropriately and results in slower generation, training, and inference.
3.2.1. Goal Encoding within the State Embedding
While the encoded Kripke model—represented through the dot language (Ellson2004)—proved effective, we encountered a second challenge. Training solely on states led the regressor to learn absolute measures of distance. This measure is independent of the underlying goal and therefore generalizes only to problems with identical goal conditions, which limits the applicability of the heuristics.
To mitigate this limitation, we extend the state embedding to include a compact encoding of the goal. This enables the regressor to learn dependencies not only between the e-state and its distance to the goal, but also between structural features of the goal itself. The design objective was to preserve a uniform input dimensionality while introducing minimal overhead. Given that the conversion process is not particularly relevant, we omit its details here and refer the reader to Appendix C for the description of the procedure.
In the resulting formulation, two types of nodes are encoded for each e-state embedding: (i) state nodes, corresponding to worlds in the epistemic state; (ii) goal nodes, corresponding to symbols describing the goal conditions. This is exemplified in Figure 1, where the left portion of the expanded hashed-based e-state (in teal) shows the goal encoding, and the right portion (in magenta) depicts the e-state itself. The two are connected via a shared graph structure using nodes and edges that use constants and special identifiers throughout the process.
In the hashed- and in the mapping-based representations, the state component remains unchanged, while goal nodes are associated with unique integer identifiers (assigned to the first occurrences of each goal symbol). In the bitmask representation, both node types share the same binary vector structure: where reflects the bitmask embedding presented above and is the binary segment reserved for goal encoding. During the experimental evaluation, we fixed the total size of to 64 bits, with , , and .
For state nodes, the goal segment is not used (), serving purely as padding to preserve dimensional uniformity across the GNN input space. For goal nodes, the segment encodes the binary representation of the integer identifier associated with the corresponding goal symbol while . Each goal symbol is uniquely represented, while agent-related components reuse the same integer indices as in the e-state encoding.
3.3. Building the Dataset
Having identified the type of data required for training, the next challenge lies in generating such data. To address this, we equipped deep with a dataset generation mode that produces pairs of epistemic states and their distances to the nearest goal. This allows training data to be collected from a small set of problems, which is then used to train GNN-based neural regressors.
The generation process works as follows: given a MEP problem, the planner performs a depth-limited Depth-First Search (DFS) to explore the reachable state space up to a specified depth (left part of Figure 1). During this traversal, all reachable goal states are identified. deep then backtracks from each goal, assigning to each epistemic state the distance to the closest goal—yielding a dataset that approximates the perfect heuristics. States from which no goal is reachable within depth are labeled with a special value.
Although conceptually simple, this process suffers from combinatorial explosion: with only 10 actions and depth 25—an overestimate of typical plan length in standard MEP benchmarks—the search space can reach nodes, making exhaustive exploration infeasible. To mitigate this, we draw on ideas from local search (Gerevini_2003), sampling subregions to maximize coverage. Our DFS incorporates probabilistic branch pruning (adaptive to depth and node count), randomized action ordering to avoid prefix bias, a node expansion cap, and duplicate e-state checks. These mechanisms enable diverse yet tractable exploration, allowing informative datasets to be built within minutes per instance.
While data generation and training are handled offline in this work, the structure of our learning pipeline naturally supports an online setup. By slightly extending the search process, the planner could incrementally collect training pairs and update the GNN once a sufficient number of samples is accumulated. Thanks to deep’s multithreading support—introduced to emulate the portfolio behavior of -EFP (DBLP:conf/prima/FabianoPSP24)—this online learning loop could run in parallel, allowing the planner to adapt and improve over time, similar to the cognitive architecture presented in 10.1145/3715709. We leave the exploration of this online learning paradigm to future work.
3.4. Training Neural Distance Estimator
Our objective is to train an estimator that, given an e-state, predicts the distance to a goal state. To this end, three design choices were made:
Discard unreachable nodes. Distance to the goal is meaningful only for e-states from which the goal can be reached. Keeping unreachable e-states in the training set injects spurious labels that raise the estimator’s variance without lowering its bias.
Limit the number of samples for any distance class. Raw rollouts produce a strongly skewed distribution, with many more short distance e-states than long distance ones, which can bias the estimator toward the majority class. We therefore limit each distance bin (in percentage) to at most of the dataset. This balance step (a) reduces variance arising from class imbalance while keeping bias low; (b) forces the network to allocate capacity uniformly throughout the distance spectrum, lowering worst‑case error and improving robustness.
Linearly normalize the distance target. To stabilize training, we linearly normalize the true distance from to , with . Let then the normalized target is In inference, we recover the original scale through
[…]
Our neural regressor is implemented in PyTorch imambi2021pytorch, and its graph encoder is built using PyTorch Geometric bielak2022pytorch. For each epistemic state (e-state), represented as a graph with nodes and edges , we construct a corresponding data object that encodes the graph structure. This object serves as input to the GNN, which processes it to produce a latent embedding. Specifically, this object includes the following:
Node identifiers: , represented as bit-vectors, in which each node is expressed as a binary vector , where is the bit-length (here ). The resulting node feature matrix is ;
Edge indices: , encoded as an index tensor , whose columns list each source–target pair;
Edge attributes: , stacked into an attribute tensor and normalized to .
The estimator is trained by minimizing the Mean Squared Error () between its predicted scalar and the true normalized distance
using the AdamW optimizer with the parameters listed in Appendix B, and where is the mini-batch size (number of samples per gradient update).
The forward pass unfolds in four stages:
Bitwise embedding of node identifiers and edge attributes: Each node bit-vector is projected through a two-layer MLP, producing dense embeddings of dimension . Similarly, each edge attribute is processed through yielding -dimensional embeddings. This bitwise encoding provides a compact and lossless representation of node identifiers while supporting continuous optimization and stable gradients.
Relational message passing with GINEConv: Two successive GINEConv (DBLP:conf/iclr/HuLGZLPL20) layers, each followed by a ReLU activation, allow each node to aggregate information from its neighbors while explicitly incorporating edge indices and attributes.
Graph-level summarization: The node embeddings from the final GINE layer are aggregated into a fixed-size representation via mean pooling: This operation is permutation-invariant and scale-independent, ensuring stable representations across graphs of different sizes. In contrast, sum pooling scales with , potentially biasing larger graphs, while attention pooling introduces learnable weights but increases variance and computational cost. Mean pooling therefore offers the best generalization–variance trade-off for distance regression.
Deep residual regression head with bounded output: The pooled graph embedding is processed by a deep residual regressor: where denotes a sequence of ResidualBlocks (Linear BatchNorm ReLU Dropout Linear skip). The sigmoid bounds the output in , and is finally clamped to to prevent numerical instabilities near the boundaries.
Overall, the combination of bitwise node embeddings, GINEConv message passing, and mean pooling yields a compact and expressive model that generalizes across variable-sized epistemic graphs while remaining numerically stable during training. The full forward-pass is detailed in Appendix E.
4. Experiments and Evaluation
To conduct our experiments, we developed deep, a modernized re-implementation of the epistemic planner EFP (icaps20; DBLP:conf/prima/FabianoPSP24). deep serves as the primary platform for our evaluation and features a modular design that supports the integration of diverse heuristics, including GNN–based ones. Full implementation details and usage instructions are available at https://github.com/FrancescoFabiano/deep.
Here we present a comparative analysis of our primary contribution, deep equipped with HFS∗—a depth-augmented variant of HFS detailed in Section 4.1—alongside GNN-based heuristics (denoted as GNN), and deep using Breadth-First Search (denoted as BFS). All executions employed bisimulation-based state reduction and visited-state checks, as introduced by icaps20. For completeness, we also provide comparisons with the different heuristics implemented in -EFP (DBLP:conf/prima/FabianoPSP24). All of these heuristics remain available in deep and can be combined with the GNN-based one in a portfolio approach. The solver is already equipped to perform this integration easily as a runtime option. We omit the results of such integration, as the focus of this paper is on the use of ML in epistemic planning, rather than on the planner itself.
While several other MEP solvers exist (muise2015planning; icaps20; KR2021-12), we focus our evaluation on BFS to emphasize the impact of incorporating learned heuristics guidance into MEP solving. All experiments were conducted with a timeout of 600 seconds on a 13th-generation Intel® Core™ i9-13900H CPU with 64 GB of system RAM and an NVIDIA RTX 4070 GPU with 8 GB of VRAM.
The evaluation encompasses several standard benchmarks in the MEP setting. For brevity, the definitions of the parameters used for training and the descriptions of the domains, are provided in Appendix B and D, respectively.
Experimental Setup.
To evaluate our contribution, we conducted experiments on standard benchmark domains. We tested two configurations: in the first, each domain had its own model trained solely on data from that domain; in the second, we evaluated the knowledge transfer capabilities of models trained on data pooled from multiple domains and tested on both seen and unseen domains.
We report a subset of experiments to highlight key trends and the impact of our contribution, while full results are provided in Appendix G.
Metrics.
Our primary evaluation metric is the number of nodes expanded during the search (), which reflects the informativeness of the heuristics. We focus on this measure because GNN currently does not leverage batch computation, making runtime comparisons less meaningful. For completeness, we also report plan length () and solving times in ms ().
For aggregate metrics, we report the Interquartile Mean () and IQR standard deviation (-) (DBLP:conf/nips/AgarwalSCCB21), focusing on problems solved by all approaches.
4.1. E-State Representation and Search Strategy
In this section, we analyze how the e-state representation and the search strategy affect the overall performance of our approach.
Specifically, we compare the three e-state representations introduced in Section 3.2 and the two search strategies, HFS and its variant HFS∗. The latter, inspired by the classical A∗ algorithm444Since the GNN-based heuristics is statistical in nature, its admissibility cannot be guaranteed; hence we do not refer to it as A∗. (FOEAD2021507; 4082128), augments the heuristics value predicted by the GNN with the search depth to form the total score:
We evaluated all six possible combinations, namely the three representations, each paired with both HFS and HFS∗. These experiments use GNN models trained on the same domain used during inference. We report aggregate results over both the Training and Test sets. We refer the reader to Appendix F for complete results. We remind that the aggregates are computed only on instances solved by all the approaches.
Table 1 reports results for the three e-state representations—map, hash, and mask—corresponding respectively to the symbolic ID, hashing, and bitmask formulations. As shown in Table 1, hash achieves the best overall performance. This can be attributed to the intrinsic richness of the e-state representation, which enables effective information retrieval even when part of the original information is discarded. Although the bitmask representation is only slightly outperformed by the hashed one in specific cases, we consider the trade-off offered by hash—lighter training requirements, faster data generation, and reduced inference time—well justifies choosing this as the best encoding.
| Solved Inst. | [ms] | |||
| map | 74/79 (93.67%) | 87 367 | 2358 9251 | 6 3 |
| hash | 75/79 (94.94%) | 45 227 | 871 3225 | 6 3 |
| mask | 75/79 (94.94%) | 55 224 | 1121 4151 | 6 3 |
Next, in Table 2, we illustrate the advantage of using HFS∗ over HFS. The results clearly indicate that HFS∗ achieves better performance, solving approximately more instances.
| Solved Inst. | [ms] | |||
| HFS | 49/79 (62.03%) | 18 27 | 1290 5528 | 8 8 |
| HFS∗ | 75/79 (94.94%) | 17 42 | 584 2276 | 6 4 |
As mentioned, for the sake of space, we presented only aggregate results. We note, however, that the trends shown in Tables 1 and 2 were consistently observed across all individual experiments. From this point onward, we therefore adopt hash combined with HFS∗ throughout the paper as our GNN configuration.
4.2. Experimental Results
[…]
| GNN | BFS | |||||||
| Solved Inst. | [ms] | Solved Inst. | [ms] | |||||
| AL | 6/7 (85.71%) | 10 0 | 274 3142 | 5 0 | 6/7 (85.71%) | 14 0 | 82 3851 | 5 0 |
| CC | 18/18 (100.00%) | 65 250 | 996 2006 | 7 4 | 18/18 (100.00%) | 610 1607 | 1610 6323 | 5 3 |
| CB | 3/3 (100.00%) | 75 860 | 279 3509 | 5 2 | 3/3 (100.00%) | 102 1260 | 118 1603 | 5 2 |
| GR | 8/12 (66.67%) | 157 380 | 6512 14016 | 6 2 | 10/12 (83.33%) | 448 2068 | 5094 12086 | 4 2 |
| SC | 19/20 (95.00%) | 49 357 | 157 511 | 10 6 | 19/20 (95.00%) | 114 372 | 131 572 | 8 4 |
| SR | 5/6 (83.33%) | 6188 17295 | 38111 106636 | 8 10 | 5/6 (83.33%) | 7918 22608 | 42146 120182 | 8 10 |
| All | 59/66 (89.39%) | 64 296 | 1001 4635 | 7 4 | 61/66 (92.42%) | 242 1080 | 922 4870 | 6 4 |
In Table 4, as a study on the knowledge transfer analysis, we compare GNN with -, a model trained using the training instances of two domains, namely CC and GR.
Finally, Table 5 presents aggregated results comparing our approach with existing heuristics in -EFP. For this evaluation, we use -, as it represents the most general-purpose configuration. A detailed comparison with individual heuristics (C_PG, L_PG, S_PG, and SUB) is provided in Appendix G. We note that the only metric with significant interpretability is the number of instances solved, since the -EFP heuristics solve a highly diverse set of instances, making other aggregated measures less informative.
| Solved Inst. | [ms] | |||
| BFS | 55/59 (93.22%) | 389 1410 | 1384 6943 | 6 4 |
| - | 55/59 (93.22%) | 288 1244 | 4116 13644 | 6 4 |
| Approach | # Solved | % Solved |
| GNN | 64/75 | |
| C_PG | 37/75 | |
| L_PG | 54/75 | |
| S_PG | 62/75 | |
| SUB | 58/75 |
4.3. Discussion
GNN demonstrates informative and robust performance across a range of planning domains, as evidenced by the aggregate metrics reported in Table 3. With the exception of the GR domain, GNN consistently reduces the number of explored nodes compared to uninformed search, highlighting the efficiency gains enabled by the learned heuristic.
The GR domain presents a particular challenge due to its sparse solution density: even small heuristic inaccuracies can result in poor search guidance, and instances with superficially similar states or goals may require substantially different plans. This variability can mislead GNN and explains its reduced effectiveness in this domain. Importantly, this observation motivates the portfolio-based heuristic selection strategy adopted in deep, where complementary heuristics can compensate when individual ones underperform. A tighter integration of GNN-based heuristics into this portfolio framework is a natural direction for future work.
[…]
Results for are comparable across methods, indicating that GNN preserves near-optimal solution quality. Solving times are likewise similar, although GNN incurs a modest overhead due to the absence of batch computation. A detailed analysis of how batch inference would affect GNN—bringing its per-node overhead closer to that of BFS—is presented in Appendix H.
Table 5 further compares GNN against individual heuristics within -EFP. GNN performs on par with the best individual heuristics, demonstrating its strength and viability as an alternative approach. At the same time, we view GNN-based heuristics as a complementary tool to enhance solver coverage, for example through integration into portfolio-based approaches, rather than as a replacement for existing heuristics.
Overall, GNN achieves consistent reductions in explored nodes and represents a scalable, learning-based step toward effective heuristics for multi-agent epistemic planning, a setting in which heuristic guidance is currently limited.
5. Limitations and Future Directions
While our approach achieves promising experimental results, several limitations remain. First, we acknowledge that our current implementation does not yet achieve competitive runtime performance compared to existing heuristics methods. This limitation is largely due to engineering considerations. A proper integration of CUDA-based computation and batch processing, combined with a search strategy capable of exploiting these features while minimizing memory exchange, would, in fact, significantly improve inference speed as detailed in Appendix H. Although we recognize this shortcoming, we emphasize that this work is foundational. Addressing the engineering challenges required for an optimized implementation would constitute a substantial effort in its own right—worthy of dedicated study—and represents an important avenue for future research. Nonetheless, our primary goal here is to provide a proof of concept highlighting the potential of heuristics learning in MEP. This is also why we focus on the number of expanded nodes as our primary evaluation metric.
Certain domains, such as AL and GR, also present unique challenges. In AL, problem instances differ only in the nesting depth of belief formulas, which results in weak learning signals. Results for this domain are therefore not very informative as GNN and BFS perform almost identically. In GR, the sparsity of valid plans limits the effectiveness of data-driven learning, exposing a current limitation of our data generation pipeline as discussed above.
As mentioned, our current GNN implementation lacks batch inference during planning, which contributes to slower runtime. While enabling batch computation is primarily an engineering task, it also raises design questions regarding when and how to accumulate batches of states for scoring. For instance, the planner could rely on BFS or alternate with other heuristics until a sufficient number of candidate states are available for batched GNN evaluation. We leave the systematic study of these strategies for future work.
Finally, an important next step is to integrate our GNN-based heuristics estimates into more advanced and potentially more effective search frameworks, such as MCTS sutton1998reinforcement. We believe this direction holds substantial promise for improving the scalability and adaptability of MEP solvers.
6. Related Works
Machine Learning in Planning.
Traditionally, planning heuristics are either hand-crafted or derived from structural features of the search space (modernApproach, Ch. 11). ML-based heuristics offer a scalable alternative, learning meaningful patterns from data (jimenez2012review; chen2024return). This is exemplified by systems like AlphaGo (silver2016mastering), where learned guidance enables scalable MCTS. Our work builds on these ideas but targets a more structured setting, where planning states are represented as Kripke structures. This introduces challenges, which we address using GNNs to extract semantic features from epistemic states. GNNs have also proven effective in classical planning, where they model relational graphs (silver2021planning), learn numeric heuristics (borelli2025learning), or guide adaptive search (du2025fast).
Recent efforts have also explored using Large Language Models (LLMs) in planning. While LLMs are ineffective as standalone planners (kambhampati2024llmscantplanhelp; DBLP:conf/icaps/PallaganiM0FLMS24), they can aid heuristics generation (kambhampati2024llmscantplanhelp; correa2025classical) or domain formalization (tantakoun-etal-2025-llms). However, due to the structured nature of MEP, we believe GNNs are a more suitable choice. We leave the integration of LLMs in this context to future work.
Multi-Agent Epistemic Planning.
Most work on MEP has focused on foundational problems such as the investigation of DEL fragments (DBLP:conf/atal/BuriganaFMT23), the definition of action languages (muise2015planning; baral2021action; DBLP:conf/aiia/BuriganaF22a), and the development of underlying representations (icaps20; DBLP:conf/jelia/BuriganaFM23). While these are fundamental contributions, this paper pursues a different goal: enabling efficient exploration via informed search.
To the best of our knowledge, only one line of work addresses this challenge, namely le2018efp; DBLP:conf/prima/FabianoPSP24, which derive heuristics using the planning graph structure. Our approach differs by employing data-driven methods. As mentioned, a future direction is to investigate the interplay between these two types of heuristics either in parallel or in conjunction.
Other recent efforts integrate ML and RL with epistemic planning but simplify epistemic state representations. engesser2025simple decompose e-states into bounded feature vectors, bridging epistemic logic and reinforcement learning. Similarly, nunn2024logic use generative models to reason over individual formulae rather than full Kripke structures, enabling different but interesting capabilities.
7. Conclusions
This work introduces a novel, learning-based approach to heuristics generation for multi-agent epistemic planning, leveraging Graph Neural Networks to guide informed search. By embedding Kripke structures and training a GNN to approximate the perfect heuristics, we enable scalable MEP planning through learning.
We investigated three distinct embeddings of Kripke structures and evaluated their effect on the heuristics accuracy. Through a comprehensive benchmarking, we identified the embedding configurations that most effectively balance expressiveness and scalability.
Our implementation, deep, demonstrates solid performance across standard benchmarks, reducing the number of explored nodes compared to uninformed search. The method also generalizes well to unseen domains and is competitive against existing heuristics.
These results highlight the potential of heuristics learning in MEP, where heuristics are scarce.
For these reasons, this work represents a foundational step toward exploiting machine learning in the context of multi-agent epistemic planning.
This research was partially funded by the EPSRC grant EP/Y028872/1, Mathematical Foundations of Intelligence: An “Erlangen Programme” for AI.
Appendix A Computational Resource Used
All experiments were performed on a Intel(R) Core(TM) i9-13900H with 20 CPU cores and 64 GB of system RAM, alongside an NVIDIA RTX 4070 GPU with 8 GB of dedicated VRAM. Note that the experiments do not need multiple repetitons as the inference and the planning process does not depend on randomness.
Overall, for the final results, sample generation, model training and inference across all experimental batches required approximately 180 hours of high‑performance computing (HPC) time. Since the project began, however, simulation trials have consumed nearly 400 hours.
Illustrative timing breakdown for a single model:
Sample generation: Approximately 15 minutes per experimental setup.
Data‑loader construction: 15 minutes for fewer than 10k samples; up to 2 hours for more than 100k samples.
Model training: 10 minutes for very small data‑loaders; up to 3 hours for very large ones.
Appendix B Parameters and Hyperparameters
In this section, we summarize all model and optimizer settings used in our experiments. Table 6 lists the architectural parameters of the --; while Table 7 gives the key AdamW optimizer hyperparameters.
| Parameter Description | Value |
| Size of hidden feature dimension in GINEConvs and regressor input. | 128 |
| Dimension of initial node embeddings (ID MLP output). | 64 |
| Dimension of initial edge embeddings (edge MLP output). | 32 |
| Hidden size in the residual regressor head. | 128 |
| Number of ResidualBlock layers in the regressor. | 3 |
| Dropout probability inside each ResidualBlock. | 0.2 |
| Lower clamp bound on final sigmoid output. | 1e-3 |
| Upper clamp bound on final sigmoid output. | 1 - 1e-3 |
| Hyperparameter Description | Value |
| Learning rate for parameter updates. | |
| Exponential decay rates for the first and second moment estimates. | (0.9, 0.999) |
| Term added to the denominator to improve numerical stability. | |
| Coefficient for decoupled weight decay (L2 regularization). | |
| Flag to enable the AMSGrad variant of AdamW. | False |
B.1. Case studies
Here, an overview of the experimental configuration parameters is provided.
| Parameter | Description | Value |
| Maximum goal distance considered | 50 | |
| Length of the input bitmask vector | 64 | |
| Lower bound for linear normalization | ||
| Upper bound for linear normalization | ||
| Maximum class imbalance threshold in the target distribution | 0.5 |
Appendix C Goal Encoding into e-State Embedding
[…]
Moreover, as shown in bolander2015complexity, the computational complexity of generating such structures is exponential in complexity with respect to the size of the formula.
To address this, we developed a custom syntax tree that parses the problem’s goal into an ad-hoc graph-based representation. This representation highlights key components such as belief operators and logical connectives, while maintaining consistent naming conventions with the state representation. Specifically, the only shared identifiers between the goal and the e-state are those that denote the same underlying entity—in our case, the agent IDs.
The procedure is presented in Algorithm 1
Appendix D Domains
Here we present the complete description of the benchmarks used to evaluate our contributions. These have been collected from the literature (kominis2015beliefs; huang2017general; cooper:hal-02147986; icaps20).
Assembly Line (AL). This domain involves two agents, each responsible for processing a separate part of a product. Each processing step may fail, and agents can inform one another of their task’s outcome. Based on this shared knowledge, the agents decide whether to assemble the product or restart. The goal is fixed—assembling the product—but the complexity varies depending on the depth of the belief formulas used in the executability conditions.
Collaboration and Communication (CC). In this domain, agents move along a corridor with rooms, in which boxes can be located. Whenever an agent enters a room, she can observe whether a specific box is present. Additionally, agents can communicate information about the boxes’ positions to other attentive agents. The goals involve both agents’ physical positions and their beliefs about the boxes.
Coin in the Box (CB). Here, agents are in a room with a locked box containing a coin, which lies either heads or tails up. Initially, no agent knows the coin’s orientation. One agent holds the key to open the box. Typical goals involve some agents learning the coin’s status, while others may need to know that someone else knows it—or remain ignorant of this fact entirely.
Grapevine (GR). In this setting, agents are distributed across rooms. Agents can freely move between rooms and share ”secrets” with any other agents present in the same room. This domain supports diverse goal types, ranging from secret sharing to creating misconceptions about others’ beliefs.
Selective Communication (SC). This domain features agents, each starting in one of the rooms arranged along a corridor. An agent may broadcast information, which is heard by all agents in the same or adjacent rooms. Each agent can move between neighboring rooms. The goals often require certain agents to know specific facts while ensuring that others remain unaware of them.
Selective Communication Enriched (SCR). This domain is a replica of SC but enriched with extra, useless actions, to decrease the density of the solutions in the search space.
Epistemic Gossip (EG – Only used in experiments presented in the Appendix). This domain extends the classic gossip problem, where information (“secrets”) must be disseminated among agents using the minimum number of calls. Unlike the traditional formulation, which requires some agents to know some secrets (epistemic depth 1), we consider arbitrary epistemic depths. For example, goals may require that all agents know that all agents know all secrets (depth 2), and so on.
Appendix E Forward Pass Neural Distance Regressor
state_graph
: node bit-vectors ( bits per node, typically );
: edge indices in COO format (Coordinate format), where each column identifies a directed edge ;
: edge attributes (scalar per edge);
: batch assignment of each node to its graph.
Notation: = bit length of node identifiers, = node embedding dimension, = edge embedding dimension, = hidden dimension in GINE layers, and = number of graphs in the batch.
Appendix F E-State Representation and Search Strategy
In what follows, we present the results for all the experiments of the e-state representation and search strategy study.
We will use the following abbreviations:
mask: e-state representation that uses symbolic ID-based embedding.
hash: e-state representation that uses hashing-based embedding.
map: e-state representation that uses bitmask-based embedding.
HFS: Best-First Search, which we abbreviate as HFS (for Heuristics-First Search) to distinguish it from BFS
HFS∗.: variant of HFS that augments the heuristics value predicted by the GNN with the search depth to form the total score:
: the number of nodes expanded during search, reflecting the informativeness of the heuristics.
: the length of the plan found.
: the solving time (in milliseconds).
: Interquartile Mean, used as an aggregate performance metric.
-: Interquartile Range, reported as a measure of variability.
: arithmetic mean, reported as a baseline aggregate metric.
: standard deviation, used to quantify variability in the data.
: indicates that the aggregate value is computed over all instances solved by that approach.
: indicates that the aggregate value is computed only over instances solved by all approaches in the comparison.
: indicates a timeout (after 600 seconds).
: indicates a missing value due to the problem not being solved within the timeout.
F.1. e-State Representations Comparison
| Instance Name | mask | hash | map | ||||||
| [ms] | [ms] | [ms] | |||||||
| Assemble_B3-pl_5 | 5 | 14 | 164 | 5 | 14 | 154 | 5 | 14 | 308 |
| Assemble_B5-pl_5 | 5 | 14 | 223 | 5 | 14 | 371 | 5 | 14 | 475 |
| Assemble_B7-pl_5 | 5 | 14 | 4368 | 5 | 14 | 4790 | 5 | 14 | 9894 |
| CC_2_2_3-pl_4 | 4 | 13 | 183 | 5 | 9 | 141 | 4 | 14 | 171 |
| CC_2_2_3-pl_6 | 6 | 62 | 424 | 6 | 10 | 106 | 6 | 178 | 1532 |
| CC_2_2_4-pl_5 | 7 | 43 | 938 | 9 | 77 | 2507 | 5 | 94 | 3483 |
| CC_2_3_4-pl_3 | 5 | 6 | 2517 | 3 | 3 | 1381 | 3 | 19 | 6067 |
| CC_2_3_4-pl_7 | 11 | 2054 | 88913 | 9 | 22 | 3244 | 7 | 4153 | 585278 |
| CC_3_2_3-pl_4 | 5 | 13 | 219 | 5 | 7 | 160 | 4 | 47 | 689 |
| CC_3_2_3-pl_5 | 6 | 51 | 888 | 6 | 45 | 463 | 5 | 136 | 2088 |
| CC_3_3_3-pl_4 | 6 | 19 | 1730 | 4 | 4 | 495 | 4 | 39 | 1337 |
| CoinBox-pl_3 | 3 | 17 | 220 | 3 | 17 | 129 | 3 | 17 | 228 |
| CoinBox-pl_6 | 6 | 381 | 3534 | 6 | 381 | 2748 | 6 | 381 | 3760 |
| SC_4_1-pl_3 | 3 | 4 | 51 | 3 | 4 | 52 | 3 | 4 | 96 |
| SC_4_2-pl_7 | 7 | 120 | 609 | 7 | 120 | 545 | 7 | 97 | 533 |
| SC_4_3-pl_5 | 5 | 17 | 116 | 5 | 17 | 62 | 5 | 17 | 98 |
| SC_8_10-pl_6 | 6 | 49 | 315 | 6 | 49 | 343 | 6 | 113 | 882 |
| SC_9_11-pl_6 | 6 | 26 | 179 | 6 | 26 | 161 | 6 | 26 | 198 |
| SC_9_11-pl_8 | 8 | 66 | 388 | 8 | 66 | 271 | 8 | 66 | 500 |
| SC_10_8-pl_10 | 10 | 10 | 174 | 10 | 10 | 97 | 10 | 10 | 108 |
| SC_10_8-pl_15 | 15 | 130 | 824 | 15 | 130 | 543 | 15 | 130 | 822 |
| SC_10_10-pl_6 | 6 | 309 | 8676 | 6 | 309 | 5998 | 6 | 298 | 8135 |
| SC_10_10-pl_7 | 7 | 647 | 17999 | 7 | 647 | 11495 | 7 | 589 | 16064 |
| SC_10_10-pl_10 | 10 | 3096 | 80030 | 10 | 3096 | 58382 | 10 | 1977 | 51904 |
| SC_10_10-pl_13 | 13 | 51 | 431 | 13 | 51 | 360 | 13 | 51 | 358 |
| () | 7 3 | 289 705 | 8565 22733 | 7 3 | 206 608 | 3800 11431 | 6 3 | 340 871 | 27800 114266 |
| - () | 6 2 | 40 106 | 748 2298 | 6 3 | 31 67 | 596 2353 | 6 2 | 66 119 | 1290 3452 |
| () | 7 3 | 289 705 | 8565 22733 | 7 3 | 206 608 | 3800 11431 | 6 3 | 340 871 | 27800 114266 |
| - () | 6 2 | 40 106 | 748 2298 | 6 3 | 31 67 | 596 2353 | 6 2 | 66 119 | 1290 3452 |
| Solved Instances | 25/25 (100.00%) | 25/25 (100.00%) | 25/25 (100.00%) | ||||||
| Instance Name | mask | hash | map | ||||||
| [ms] | [ms] | [ms] | |||||||
| Assemble_B10-pl_5 | |||||||||
| Assemble_B2-pl_5 | 5 | 14 | 264 | 5 | 14 | 264 | 5 | 14 | 285 |
| Assemble_B4-pl_5 | 5 | 14 | 275 | 5 | 14 | 332 | 5 | 14 | 330 |
| Assemble_B6-pl_5 | 5 | 14 | 804 | 5 | 14 | 826 | 5 | 14 | 1127 |
| Assemble_B8-pl_5 | 5 | 14 | 24655 | 5 | 14 | 25636 | 5 | 14 | 48340 |
| Assemble_B9-pl_5 | 5 | 14 | 295842 | 5 | 14 | 296625 | |||
| Assemble_C-pl_5 | 5 | 14 | 124 | 5 | 14 | 237 | 5 | 14 | 238 |
| CC_2_2_3-pl_3 | 3 | 9 | 111 | 4 | 11 | 108 | 3 | 9 | 181 |
| CC_2_2_3-pl_5 | 5 | 17 | 272 | 5 | 9 | 134 | 5 | 77 | 684 |
| CC_2_2_3-pl_7 | 7 | 163 | 2390 | 8 | 33837 | 214811 | 7 | 558 | 4358 |
| CC_2_2_3-pl_8 | 8 | 1327 | 18235 | 8 | 247 | 2434 | 8 | 2539 | 27312 |
| CC_2_2_4-pl_3 | 3 | 3 | 217 | 3 | 3 | 279 | 3 | 11 | 373 |
| CC_2_2_4-pl_4 | 6 | 7 | 324 | 6 | 9 | 212 | 4 | 23 | 778 |
| CC_2_2_4-pl_6 | 6 | 21 | 1251 | 7 | 26 | 296 | 6 | 363 | 9536 |
| CC_2_2_4-pl_7 | 8 | 129 | 3989 | 8 | 498 | 4701 | 7 | 2084 | 45668 |
| CC_2_3_4-pl_4 | 5 | 5 | 1995 | 6 | 10 | 1655 | 4 | 42 | 12130 |
| CC_2_3_4-pl_5 | 10 | 132 | 20181 | 12 | 61 | 5848 | 5 | 127 | 45028 |
| CC_2_3_4-pl_6 | 11 | 1269 | 65775 | 8 | 11 | 3383 | 6 | 675 | 124726 |
| CC_3_2_3-pl_3 | 3 | 4 | 228 | 3 | 3 | 196 | 3 | 14 | 318 |
| CC_3_2_3-pl_6 | 6 | 74 | 1201 | 6 | 113 | 1092 | 6 | 367 | 6443 |
| CC_3_2_3-pl_7 | 8 | 185 | 3058 | 7 | 1835 | 24464 | 7 | 1548 | 26110 |
| CC_3_3_3-pl_3 | 3 | 3 | 218 | 3 | 4 | 367 | 3 | 15 | 790 |
| CC_3_3_3-pl_5 | 6 | 26 | 1975 | 7 | 22 | 534 | 5 | 253 | 8316 |
| CC_3_3_3-pl_6 | 8 | 226 | 13395 | 6 | 1942 | 73680 | |||
| CC_3_3_3-pl_7 | 8 | 1476 | 59481 | 8 | 79 | 1255 | 7 | 11726 | 405444 |
| CoinBox-pl_2 | 2 | 2 | 52 | 2 | 2 | 50 | 2 | 2 | 109 |
| CoinBox-pl_5 | 5 | 77 | 425 | 5 | 77 | 552 | 5 | 78 | 728 |
| CoinBox-pl_7 | 7 | 1817 | 10395 | 7 | 1817 | 13110 | 7 | 1793 | 17758 |
| SC_4_1-pl_5 | 5 | 17 | 125 | 5 | 17 | 99 | 5 | 17 | 93 |
| SC_4_2-pl_5 | 5 | 15 | 132 | 5 | 15 | 99 | 5 | 30 | 184 |
| SC_4_2-pl_8 | 8 | 339 | 1639 | 8 | 339 | 1316 | 8 | 184 | 923 |
| SC_4_3-pl_6 | 6 | 21 | 162 | 6 | 21 | 113 | 6 | 21 | 93 |
| SC_4_3-pl_8 | 8 | 59 | 186 | 8 | 59 | 156 | 8 | 59 | 177 |
| SC_4_4-pl_5 | 5 | 17 | 125 | 5 | 17 | 87 | 5 | 17 | 91 |
| SC_8_10-pl_8 | 8 | 542 | 5157 | 8 | 542 | 3841 | 8 | 485 | 4121 |
| SC_8_10-pl_9 | 9 | 1660 | 17268 | 9 | 1660 | 14557 | 9 | 1772 | 16936 |
| SC_8_10-pl_12 | 12 | 31662 | 199243 | 12 | 31662 | 203157 | 12 | 35063 | 281914 |
| SC_9_11-pl_4 | 4 | 8 | 143 | 4 | 8 | 141 | 4 | 8 | 139 |
| SC_9_11-pl_5 | 5 | 11 | 152 | 5 | 11 | 96 | 5 | 11 | 122 |
| SC_9_11-pl_7 | 7 | 41 | 431 | 7 | 41 | 299 | 7 | 41 | 326 |
| SC_9_11-pl_9 | 9 | 238 | 1895 | 9 | 238 | 2158 | 9 | 237 | 2041 |
| SC_9_11-pl_10 | 10 | 663 | 6151 | 10 | 663 | 5467 | 10 | 657 | 6013 |
| SC_9_11-pl_11 | 11 | 1262 | 15073 | 11 | 1262 | 10960 | 11 | 1272 | 12372 |
| SC_10_8-pl_9 | |||||||||
| SC_10_8-pl_14 | 14 | 93 | 481 | 14 | 93 | 378 | 14 | 93 | 572 |
| SC_10_10-pl_2 | 2 | 6 | 280 | 2 | 6 | 173 | 2 | 6 | 311 |
| SC_10_10-pl_3 | 3 | 10 | 239 | 3 | 10 | 248 | 3 | 10 | 347 |
| SC_10_10-pl_9 | 9 | 10 | 188 | 9 | 10 | 120 | 9 | 10 | 117 |
| SC_10_10-pl_9 | 9 | 886 | 20614 | 9 | 886 | 17444 | 9 | 1248 | 32644 |
| SC_10_10-pl_10 | 10 | 15 | 191 | 10 | 15 | 158 | 10 | 15 | 161 |
| SC_10_10-pl_13 | 13 | 15826 | 285398 | 13 | 15826 | 330014 | 13 | 17305 | 472785 |
| SC_10_10-pl_14 | 14 | 30967 | 582175 | ||||||
| SC_10_10-pl_17 | 17 | 946 | 7240 | 17 | 946 | 5175 | 17 | 942 | 7546 |
| SC_10_10-pl_17 | |||||||||
| () | 7 3 | 1228 4886 | 21793 62537 | 7 3 | 2481 7827 | 35557 106042 | 7 3 | 1711 5628 | 34711 94709 |
| - () | 6 4 | 66 300 | 1758 9389 | 6 4 | 61 447 | 1335 5215 | 6 3 | 156 661 | 3908 16651 |
| () | 7 3 | 1274 4981 | 16258 49770 | 7 3 | 1939 6818 | 18730 61527 | 7 3 | 1706 5687 | 33899 95522 |
| - () | 7 4 | 73 376 | 1465 6213 | 6 3 | 63 368 | 1126 4650 | 6 3 | 137 648 | 3365 13240 |
| Solved Instances | 50/54 (92.59%) | 50/54 (92.59%) | 49/54 (90.74%) | ||||||
| Instance Name | mask | hash | map | ||||||
| [ms] | [ms] | [ms] | |||||||
| Assemble_B10-pl_5 | |||||||||
| Assemble_B2-pl_5 | 5 | 14 | 264 | 5 | 14 | 264 | 5 | 14 | 285 |
| Assemble_B3-pl_5 | 5 | 14 | 164 | 5 | 14 | 154 | 5 | 14 | 308 |
| Assemble_B4-pl_5 | 5 | 14 | 275 | 5 | 14 | 332 | 5 | 14 | 330 |
| Assemble_B5-pl_5 | 5 | 14 | 223 | 5 | 14 | 371 | 5 | 14 | 475 |
| Assemble_B6-pl_5 | 5 | 14 | 804 | 5 | 14 | 826 | 5 | 14 | 1127 |
| Assemble_B7-pl_5 | 5 | 14 | 4368 | 5 | 14 | 4790 | 5 | 14 | 9894 |
| Assemble_B8-pl_5 | 5 | 14 | 24655 | 5 | 14 | 25636 | 5 | 14 | 48340 |
| Assemble_B9-pl_5 | 5 | 14 | 295842 | 5 | 14 | 296625 | |||
| Assemble_C-pl_5 | 5 | 14 | 124 | 5 | 14 | 237 | 5 | 14 | 238 |
| CC_2_2_3-pl_3 | 3 | 9 | 111 | 4 | 11 | 108 | 3 | 9 | 181 |
| CC_2_2_3-pl_4 | 4 | 13 | 183 | 5 | 9 | 141 | 4 | 14 | 171 |
| CC_2_2_3-pl_5 | 5 | 17 | 272 | 5 | 9 | 134 | 5 | 77 | 684 |
| CC_2_2_3-pl_6 | 6 | 62 | 424 | 6 | 10 | 106 | 6 | 178 | 1532 |
| CC_2_2_3-pl_7 | 7 | 163 | 2390 | 8 | 33837 | 214811 | 7 | 558 | 4358 |
| CC_2_2_3-pl_8 | 8 | 1327 | 18235 | 8 | 247 | 2434 | 8 | 2539 | 27312 |
| CC_2_2_4-pl_3 | 3 | 3 | 217 | 3 | 3 | 279 | 3 | 11 | 373 |
| CC_2_2_4-pl_4 | 6 | 7 | 324 | 6 | 9 | 212 | 4 | 23 | 778 |
| CC_2_2_4-pl_5 | 7 | 43 | 938 | 9 | 77 | 2507 | 5 | 94 | 3483 |
| CC_2_2_4-pl_6 | 6 | 21 | 1251 | 7 | 26 | 296 | 6 | 363 | 9536 |
| CC_2_2_4-pl_7 | 8 | 129 | 3989 | 8 | 498 | 4701 | 7 | 2084 | 45668 |
| CC_2_3_4-pl_3 | 5 | 6 | 2517 | 3 | 3 | 1381 | 3 | 19 | 6067 |
| CC_2_3_4-pl_4 | 5 | 5 | 1995 | 6 | 10 | 1655 | 4 | 42 | 12130 |
| CC_2_3_4-pl_5 | 10 | 132 | 20181 | 12 | 61 | 5848 | 5 | 127 | 45028 |
| CC_2_3_4-pl_6 | 11 | 1269 | 65775 | 8 | 11 | 3383 | 6 | 675 | 124726 |
| CC_2_3_4-pl_7 | 11 | 2054 | 88913 | 9 | 22 | 3244 | 7 | 4153 | 585278 |
| CC_3_2_3-pl_3 | 3 | 4 | 228 | 3 | 3 | 196 | 3 | 14 | 318 |
| CC_3_2_3-pl_4 | 5 | 13 | 219 | 5 | 7 | 160 | 4 | 47 | 689 |
| CC_3_2_3-pl_5 | 6 | 51 | 888 | 6 | 45 | 463 | 5 | 136 | 2088 |
| CC_3_2_3-pl_6 | 6 | 74 | 1201 | 6 | 113 | 1092 | 6 | 367 | 6443 |
| CC_3_2_3-pl_7 | 8 | 185 | 3058 | 7 | 1835 | 24464 | 7 | 1548 | 26110 |
| CC_3_3_3-pl_3 | 3 | 3 | 218 | 3 | 4 | 367 | 3 | 15 | 790 |
| CC_3_3_3-pl_4 | 6 | 19 | 1730 | 4 | 4 | 495 | 4 | 39 | 1337 |
| CC_3_3_3-pl_5 | 6 | 26 | 1975 | 7 | 22 | 534 | 5 | 253 | 8316 |
| CC_3_3_3-pl_6 | 8 | 226 | 13395 | 6 | 1942 | 73680 | |||
| CC_3_3_3-pl_7 | 8 | 1476 | 59481 | 8 | 79 | 1255 | 7 | 11726 | 405444 |
| CoinBox-pl_2 | 2 | 2 | 52 | 2 | 2 | 50 | 2 | 2 | 109 |
| CoinBox-pl_3 | 3 | 17 | 220 | 3 | 17 | 129 | 3 | 17 | 228 |
| CoinBox-pl_5 | 5 | 77 | 425 | 5 | 77 | 552 | 5 | 78 | 728 |
| CoinBox-pl_6 | 6 | 381 | 3534 | 6 | 381 | 2748 | 6 | 381 | 3760 |
| CoinBox-pl_7 | 7 | 1817 | 10395 | 7 | 1817 | 13110 | 7 | 1793 | 17758 |
| SC_4_1-pl_3 | 3 | 4 | 51 | 3 | 4 | 52 | 3 | 4 | 96 |
| SC_4_1-pl_5 | 5 | 17 | 125 | 5 | 17 | 99 | 5 | 17 | 93 |
| SC_4_2-pl_5 | 5 | 15 | 132 | 5 | 15 | 99 | 5 | 30 | 184 |
| SC_4_2-pl_7 | 7 | 120 | 609 | 7 | 120 | 545 | 7 | 97 | 533 |
| SC_4_2-pl_8 | 8 | 339 | 1639 | 8 | 339 | 1316 | 8 | 184 | 923 |
| SC_4_3-pl_5 | 5 | 17 | 116 | 5 | 17 | 62 | 5 | 17 | 98 |
| SC_4_3-pl_6 | 6 | 21 | 162 | 6 | 21 | 113 | 6 | 21 | 93 |
| SC_4_3-pl_8 | 8 | 59 | 186 | 8 | 59 | 156 | 8 | 59 | 177 |
| SC_4_4-pl_5 | 5 | 17 | 125 | 5 | 17 | 87 | 5 | 17 | 91 |
| SC_8_10-pl_6 | 6 | 49 | 315 | 6 | 49 | 343 | 6 | 113 | 882 |
| SC_8_10-pl_8 | 8 | 542 | 5157 | 8 | 542 | 3841 | 8 | 485 | 4121 |
| SC_8_10-pl_9 | 9 | 1660 | 17268 | 9 | 1660 | 14557 | 9 | 1772 | 16936 |
| SC_8_10-pl_12 | 12 | 31662 | 199243 | 12 | 31662 | 203157 | 12 | 35063 | 281914 |
| SC_9_11-pl_4 | 4 | 8 | 143 | 4 | 8 | 141 | 4 | 8 | 139 |
| SC_9_11-pl_5 | 5 | 11 | 152 | 5 | 11 | 96 | 5 | 11 | 122 |
| SC_9_11-pl_6 | 6 | 26 | 179 | 6 | 26 | 161 | 6 | 26 | 198 |
| SC_9_11-pl_7 | 7 | 41 | 431 | 7 | 41 | 299 | 7 | 41 | 326 |
| SC_9_11-pl_8 | 8 | 66 | 388 | 8 | 66 | 271 | 8 | 66 | 500 |
| SC_9_11-pl_9 | 9 | 238 | 1895 | 9 | 238 | 2158 | 9 | 237 | 2041 |
| SC_9_11-pl_10 | 10 | 663 | 6151 | 10 | 663 | 5467 | 10 | 657 | 6013 |
| SC_9_11-pl_11 | 11 | 1262 | 15073 | 11 | 1262 | 10960 | 11 | 1272 | 12372 |
| SC_10_8-pl_9 | |||||||||
| SC_10_8-pl_10 | 10 | 10 | 174 | 10 | 10 | 97 | 10 | 10 | 108 |
| SC_10_8-pl_14 | 14 | 93 | 481 | 14 | 93 | 378 | 14 | 93 | 572 |
| SC_10_8-pl_15 | 15 | 130 | 824 | 15 | 130 | 543 | 15 | 130 | 822 |
| SC_10_10-pl_2 | 2 | 6 | 280 | 2 | 6 | 173 | 2 | 6 | 311 |
| SC_10_10-pl_3 | 3 | 10 | 239 | 3 | 10 | 248 | 3 | 10 | 347 |
| SC_10_10-pl_6 | 6 | 309 | 8676 | 6 | 309 | 5998 | 6 | 298 | 8135 |
| SC_10_10-pl_7 | 7 | 647 | 17999 | 7 | 647 | 11495 | 7 | 589 | 16064 |
| SC_10_10-pl_9 | 9 | 10 | 188 | 9 | 10 | 120 | 9 | 10 | 117 |
| SC_10_10-pl_9 | 9 | 886 | 20614 | 9 | 886 | 17444 | 9 | 1248 | 32644 |
| SC_10_10-pl_10 | 10 | 15 | 191 | 10 | 15 | 158 | 10 | 15 | 161 |
| SC_10_10-pl_10 | 10 | 3096 | 80030 | 10 | 3096 | 58382 | 10 | 1977 | 51904 |
| SC_10_10-pl_13 | 13 | 15826 | 285398 | 13 | 15826 | 330014 | 13 | 17305 | 472785 |
| SC_10_10-pl_13 | 13 | 51 | 431 | 13 | 51 | 360 | 13 | 51 | 358 |
| SC_10_10-pl_14 | 14 | 30967 | 582175 | ||||||
| SC_10_10-pl_17 | 17 | 946 | 7240 | 17 | 946 | 5175 | 17 | 942 | 7546 |
| SC_10_10-pl_17 | |||||||||
| () | 7 3 | 915 4034 | 17383 53089 | 7 3 | 1723 6489 | 24971 88115 | 6 3 | 1248 4653 | 32376 101790 |
| - () | 6 3 | 54 218 | 1255 5436 | 6 4 | 44 232 | 971 4112 | 6 3 | 87 445 | 2415 9514 |
| () | 7 3 | 937 4087 | 13624 42651 | 7 3 | 1346 5601 | 13617 50834 | 7 3 | 1238 4684 | 31810 102369 |
| - () | 6 3 | 55 224 | 1121 4151 | 6 3 | 45 227 | 871 3225 | 6 3 | 87 367 | 2358 9251 |
| Solved Instances | 75/79 (94.94%) | 75/79 (94.94%) | 74/79 (93.67%) | ||||||
F.2. Symbolic Mapping Representation Search Strategies Comparison
| Instance Name | HFS∗ | HFS | ||||
| [ms] | [ms] | |||||
| Assemble_B3-pl_5 | 5 | 14 | 308 | 5 | 14 | 302 |
| Assemble_B5-pl_5 | 5 | 14 | 475 | 5 | 14 | 448 |
| Assemble_B7-pl_5 | 5 | 14 | 9894 | 5 | 14 | 10020 |
| CC_2_2_3-pl_4 | 4 | 14 | 171 | 4 | 9 | 182 |
| CC_2_2_3-pl_6 | 6 | 178 | 1532 | 7 | 25 | 291 |
| CC_2_2_4-pl_5 | 5 | 94 | 3483 | 12 | 70 | 2787 |
| CC_2_3_4-pl_3 | 3 | 19 | 6067 | 24 | 199 | 49682 |
| CC_2_3_4-pl_7 | 7 | 4153 | 585278 | |||
| CC_3_2_3-pl_4 | 4 | 47 | 689 | 5 | 16 | 254 |
| CC_3_2_3-pl_5 | 5 | 136 | 2088 | 8 | 39 | 500 |
| CC_3_3_3-pl_4 | 4 | 39 | 1337 | 4 | 10 | 599 |
| CoinBox-pl_3 | 3 | 17 | 228 | 4 | 12 | 201 |
| CoinBox-pl_6 | 6 | 381 | 3760 | |||
| SC_4_1-pl_3 | 3 | 4 | 96 | 3 | 4 | 109 |
| SC_4_2-pl_7 | 7 | 97 | 533 | |||
| SC_4_3-pl_5 | 5 | 17 | 98 | |||
| SC_8_10-pl_6 | 6 | 113 | 882 | 1951 | 24489 | 384947 |
| SC_9_11-pl_6 | 6 | 26 | 198 | 509 | 5116 | 48319 |
| SC_9_11-pl_8 | 8 | 66 | 500 | 1099 | 12447 | 152051 |
| SC_10_8-pl_10 | 10 | 10 | 108 | 10 | 10 | 137 |
| SC_10_8-pl_15 | 15 | 130 | 822 | |||
| SC_10_10-pl_6 | 6 | 298 | 8135 | |||
| SC_10_10-pl_7 | 7 | 589 | 16064 | |||
| SC_10_10-pl_10 | 10 | 1977 | 51904 | |||
| SC_10_10-pl_13 | 13 | 51 | 358 | 17 | 49 | 400 |
| () | 6 3 | 340 871 | 27800 114266 | 216 513 | 2502 6299 | 38308 94330 |
| - () | 6 2 | 66 119 | 1290 3452 | 8 12 | 28 58 | 1733 9766 |
| () | 6 3 | 50 49 | 1671 2543 | 216 513 | 2502 6299 | 38308 94330 |
| - () | 5 2 | 29 52 | 701 1304 | 8 12 | 28 58 | 1733 9766 |
| Solved Instances | 25/25 (100.00%) | 17/25 (68.00%) | ||||
| Instance Name | HFS∗ | HFS | ||||
| [ms] | [ms] | |||||
| Assemble_B10-pl_5 | ||||||
| Assemble_B2-pl_5 | 5 | 14 | 285 | 5 | 14 | 287 |
| Assemble_B4-pl_5 | 5 | 14 | 330 | 5 | 14 | 357 |
| Assemble_B6-pl_5 | 5 | 14 | 1127 | 5 | 14 | 1190 |
| Assemble_B8-pl_5 | 5 | 14 | 48340 | 5 | 14 | 49503 |
| Assemble_B9-pl_5 | ||||||
| Assemble_C-pl_5 | 5 | 14 | 238 | 5 | 14 | 257 |
| CC_2_2_3-pl_3 | 3 | 9 | 181 | 4 | 10 | 184 |
| CC_2_2_3-pl_5 | 5 | 77 | 684 | 7 | 28 | 255 |
| CC_2_2_3-pl_7 | 7 | 558 | 4358 | 44 | 363 | 2249 |
| CC_2_2_3-pl_8 | 8 | 2539 | 27312 | |||
| CC_2_2_4-pl_3 | 3 | 11 | 373 | 4 | 9 | 370 |
| CC_2_2_4-pl_4 | 4 | 23 | 778 | 7 | 28 | 866 |
| CC_2_2_4-pl_6 | 6 | 363 | 9536 | 13 | 79 | 2764 |
| CC_2_2_4-pl_7 | 7 | 2084 | 45668 | |||
| CC_2_3_4-pl_4 | 4 | 42 | 12130 | 24 | 201 | 49087 |
| CC_2_3_4-pl_5 | 5 | 127 | 45028 | 13 | 85 | 27215 |
| CC_2_3_4-pl_6 | 6 | 675 | 124726 | 29 | 256 | 56262 |
| CC_3_2_3-pl_3 | 3 | 14 | 318 | 9 | 48 | 533 |
| CC_3_2_3-pl_6 | 6 | 367 | 6443 | 6 | 21 | 323 |
| CC_3_2_3-pl_7 | 7 | 1548 | 26110 | |||
| CC_3_3_3-pl_3 | 3 | 15 | 790 | 5 | 17 | 828 |
| CC_3_3_3-pl_5 | 5 | 253 | 8316 | 18 | 137 | 3073 |
| CC_3_3_3-pl_6 | 6 | 1942 | 73680 | 31 | 279 | 5961 |
| CC_3_3_3-pl_7 | 7 | 11726 | 405444 | |||
| CoinBox-pl_2 | 2 | 2 | 109 | 2 | 2 | 106 |
| CoinBox-pl_5 | 5 | 78 | 728 | 5 | 16 | 265 |
| CoinBox-pl_7 | 7 | 1793 | 17758 | |||
| SC_4_1-pl_5 | 5 | 17 | 93 | |||
| SC_4_2-pl_5 | 5 | 30 | 184 | |||
| SC_4_2-pl_8 | 8 | 184 | 923 | |||
| SC_4_3-pl_6 | 6 | 21 | 93 | |||
| SC_4_3-pl_8 | 8 | 59 | 177 | |||
| SC_4_4-pl_5 | 5 | 17 | 91 | |||
| SC_8_10-pl_8 | 8 | 485 | 4121 | |||
| SC_8_10-pl_9 | 9 | 1772 | 16936 | |||
| SC_8_10-pl_12 | 12 | 35063 | 281914 | |||
| SC_9_11-pl_4 | 4 | 8 | 139 | 6 | 13 | 150 |
| SC_9_11-pl_5 | 5 | 11 | 122 | 8 | 21 | 189 |
| SC_9_11-pl_7 | 7 | 41 | 326 | 1286 | 15107 | 202732 |
| SC_9_11-pl_9 | 9 | 237 | 2041 | |||
| SC_9_11-pl_10 | 10 | 657 | 6013 | |||
| SC_9_11-pl_11 | 11 | 1272 | 12372 | |||
| SC_10_8-pl_9 | ||||||
| SC_10_8-pl_14 | 14 | 93 | 572 | 16 | 33 | 252 |
| SC_10_10-pl_2 | 2 | 6 | 311 | 3 | 8 | 315 |
| SC_10_10-pl_3 | 3 | 10 | 347 | |||
| SC_10_10-pl_9 | 9 | 10 | 117 | 9 | 10 | 132 |
| SC_10_10-pl_9 | 9 | 1248 | 32644 | 12 | 57 | 1589 |
| SC_10_10-pl_10 | 10 | 15 | 161 | 11 | 17 | 176 |
| SC_10_10-pl_13 | 13 | 17305 | 472785 | |||
| SC_10_10-pl_14 | ||||||
| SC_10_10-pl_17 | 17 | 942 | 7546 | |||
| SC_10_10-pl_17 | ||||||
| () | 7 3 | 1711 5628 | 34711 94709 | 55 233 | 583 2746 | 14051 39043 |
| - () | 6 3 | 156 661 | 3908 16651 | 8 8 | 27 65 | 830 2509 |
| () | 5 2 | 209 424 | 12858 27539 | 55 233 | 583 2746 | 14051 39043 |
| - () | 5 2 | 40 113 | 1716 8031 | 8 8 | 27 65 | 830 2509 |
| Solved Instances | 49/54 (90.74%) | 29/54 (53.70%) | ||||
F.3. Hashed-Based Representation Search Strategies Comparison
| Instance Name | HFS∗ | HFS | ||||
| [ms] | [ms] | |||||
| Assemble_B3-pl_5 | 5 | 14 | 154 | 5 | 14 | 249 |
| Assemble_B5-pl_5 | 5 | 14 | 371 | 5 | 14 | 245 |
| Assemble_B7-pl_5 | 5 | 14 | 4790 | 5 | 14 | 4660 |
| CC_2_2_3-pl_4 | 5 | 9 | 141 | 7 | 12 | 123 |
| CC_2_2_3-pl_6 | 6 | 10 | 106 | 9 | 18 | 246 |
| CC_2_2_4-pl_5 | 9 | 77 | 2507 | 13 | 38 | 1221 |
| CC_2_3_4-pl_3 | 3 | 3 | 1381 | 3 | 3 | 918 |
| CC_2_3_4-pl_7 | 9 | 22 | 3244 | 9 | 19 | 4247 |
| CC_3_2_3-pl_4 | 5 | 7 | 160 | 6 | 8 | 315 |
| CC_3_2_3-pl_5 | 6 | 45 | 463 | |||
| CC_3_3_3-pl_4 | 4 | 4 | 495 | 4 | 4 | 424 |
| CoinBox-pl_3 | 3 | 17 | 129 | 4 | 12 | 142 |
| CoinBox-pl_6 | 6 | 381 | 2748 | |||
| SC_4_1-pl_3 | 3 | 4 | 52 | 3 | 4 | 125 |
| SC_4_2-pl_7 | 7 | 120 | 545 | |||
| SC_4_3-pl_5 | 5 | 17 | 62 | |||
| SC_8_10-pl_6 | 6 | 49 | 343 | 483 | 4776 | 58158 |
| SC_9_11-pl_6 | 6 | 26 | 161 | 509 | 5116 | 55844 |
| SC_9_11-pl_8 | 8 | 66 | 271 | 1098 | 12438 | 186999 |
| SC_10_8-pl_10 | 10 | 10 | 97 | 10 | 10 | 112 |
| SC_10_8-pl_15 | 15 | 130 | 543 | |||
| SC_10_10-pl_6 | 6 | 309 | 5998 | |||
| SC_10_10-pl_7 | 7 | 647 | 11495 | |||
| SC_10_10-pl_10 | 10 | 3096 | 58382 | |||
| SC_10_10-pl_13 | 13 | 51 | 360 | 17 | 49 | 519 |
| () | 7 3 | 206 608 | 3800 11431 | 129 289 | 1326 3197 | 18503 45821 |
| - () | 6 3 | 31 67 | 596 2353 | 8 8 | 17 28 | 932 4002 |
| () | 6 3 | 23 22 | 868 1321 | 129 289 | 1326 3197 | 18503 45821 |
| - () | 6 3 | 15 17 | 273 354 | 8 8 | 17 28 | 932 4002 |
| Solved Instances | 25/25 (100.00%) | 17/25 (68.00%) | ||||
| Instance Name | HFS∗ | HFS | ||||
| [ms] | [ms] | |||||
| Assemble_B10-pl_5 | ||||||
| Assemble_B2-pl_5 | 5 | 14 | 264 | 5 | 14 | 125 |
| Assemble_B4-pl_5 | 5 | 14 | 332 | 5 | 14 | 167 |
| Assemble_B6-pl_5 | 5 | 14 | 826 | 5 | 14 | 689 |
| Assemble_B8-pl_5 | 5 | 14 | 25636 | 5 | 14 | 25949 |
| Assemble_B9-pl_5 | 5 | 14 | 296625 | 5 | 14 | 379798 |
| Assemble_C-pl_5 | 5 | 14 | 237 | 5 | 14 | 165 |
| CC_2_2_3-pl_3 | 4 | 11 | 108 | 22 | 146 | 2225 |
| CC_2_2_3-pl_5 | 5 | 9 | 134 | 10 | 24 | 563 |
| CC_2_2_3-pl_7 | 8 | 33837 | 214811 | |||
| CC_2_2_3-pl_8 | 8 | 247 | 2434 | 11 | 36 | 763 |
| CC_2_2_4-pl_3 | 3 | 3 | 279 | 3 | 3 | 290 |
| CC_2_2_4-pl_4 | 6 | 9 | 212 | 8 | 11 | 689 |
| CC_2_2_4-pl_6 | 7 | 26 | 296 | |||
| CC_2_2_4-pl_7 | 8 | 498 | 4701 | |||
| CC_2_3_4-pl_4 | 6 | 10 | 1655 | 13 | 19 | 4652 |
| CC_2_3_4-pl_5 | 12 | 61 | 5848 | 13 | 24 | 5774 |
| CC_2_3_4-pl_6 | 8 | 11 | 3383 | 8 | 10 | 5843 |
| CC_3_2_3-pl_3 | 3 | 3 | 196 | 3 | 3 | 330 |
| CC_3_2_3-pl_6 | 6 | 113 | 1092 | |||
| CC_3_2_3-pl_7 | 7 | 1835 | 24464 | |||
| CC_3_3_3-pl_3 | 3 | 4 | 367 | 3 | 3 | 295 |
| CC_3_3_3-pl_5 | 7 | 22 | 534 | 8 | 11 | 670 |
| CC_3_3_3-pl_6 | ||||||
| CC_3_3_3-pl_7 | 8 | 79 | 1255 | |||
| CoinBox-pl_2 | 2 | 2 | 50 | 2 | 2 | 64 |
| CoinBox-pl_5 | 5 | 77 | 552 | 5 | 16 | 220 |
| CoinBox-pl_7 | 7 | 1817 | 13110 | |||
| SC_4_1-pl_5 | 5 | 17 | 99 | |||
| SC_4_2-pl_5 | 5 | 15 | 99 | |||
| SC_4_2-pl_8 | 8 | 339 | 1316 | |||
| SC_4_3-pl_6 | 6 | 21 | 113 | |||
| SC_4_3-pl_8 | 8 | 59 | 156 | |||
| SC_4_4-pl_5 | 5 | 17 | 87 | |||
| SC_8_10-pl_8 | 8 | 542 | 3841 | 1586 | 19087 | 326379 |
| SC_8_10-pl_9 | 9 | 1660 | 14557 | 1586 | 19087 | 319399 |
| SC_8_10-pl_12 | 12 | 31662 | 203157 | |||
| SC_9_11-pl_4 | 4 | 8 | 141 | 6 | 13 | 172 |
| SC_9_11-pl_5 | 5 | 11 | 96 | 8 | 21 | 296 |
| SC_9_11-pl_7 | 7 | 41 | 299 | 1286 | 15107 | 195915 |
| SC_9_11-pl_9 | 9 | 238 | 2158 | 1099 | 12457 | 150920 |
| SC_9_11-pl_10 | 10 | 663 | 5467 | 1100 | 12470 | 176047 |
| SC_9_11-pl_11 | 11 | 1262 | 10960 | 1100 | 12472 | 173729 |
| SC_10_8-pl_9 | ||||||
| SC_10_8-pl_14 | 14 | 93 | 378 | 16 | 33 | 273 |
| SC_10_10-pl_2 | 2 | 6 | 173 | 4 | 13 | 720 |
| SC_10_10-pl_3 | 3 | 10 | 248 | |||
| SC_10_10-pl_9 | 9 | 10 | 120 | 9 | 10 | 225 |
| SC_10_10-pl_9 | 9 | 886 | 17444 | 13 | 65 | 946 |
| SC_10_10-pl_10 | 10 | 15 | 158 | 11 | 17 | 273 |
| SC_10_10-pl_13 | 13 | 15826 | 330014 | |||
| SC_10_10-pl_14 | 14 | 30967 | 582175 | |||
| SC_10_10-pl_17 | 17 | 946 | 5175 | |||
| SC_10_10-pl_17 | ||||||
| () | 7 3 | 2481 7827 | 35557 106042 | 249 510 | 2851 6028 | 55455 108611 |
| - () | 6 4 | 61 447 | 1335 5215 | 8 9 | 19 31 | 1488 10596 |
| () | 7 3 | 187 391 | 12358 51387 | 249 510 | 2851 6028 | 55455 108611 |
| - () | 7 4 | 22 71 | 882 3307 | 8 9 | 19 31 | 1488 10596 |
| Solved Instances | 50/54 (92.59%) | 32/54 (59.26%) | ||||
| Instance Name | HFS∗ | HFS | ||||
| [ms] | [ms] | |||||
| Assemble_B10-pl_5 | ||||||
| Assemble_B2-pl_5 | 5 | 14 | 264 | 5 | 14 | 125 |
| Assemble_B3-pl_5 | 5 | 14 | 154 | 5 | 14 | 249 |
| Assemble_B4-pl_5 | 5 | 14 | 332 | 5 | 14 | 167 |
| Assemble_B5-pl_5 | 5 | 14 | 371 | 5 | 14 | 245 |
| Assemble_B6-pl_5 | 5 | 14 | 826 | 5 | 14 | 689 |
| Assemble_B7-pl_5 | 5 | 14 | 4790 | 5 | 14 | 4660 |
| Assemble_B8-pl_5 | 5 | 14 | 25636 | 5 | 14 | 25949 |
| Assemble_B9-pl_5 | 5 | 14 | 296625 | 5 | 14 | 379798 |
| Assemble_C-pl_5 | 5 | 14 | 237 | 5 | 14 | 165 |
| CC_2_2_3-pl_3 | 4 | 11 | 108 | 22 | 146 | 2225 |
| CC_2_2_3-pl_4 | 5 | 9 | 141 | 7 | 12 | 123 |
| CC_2_2_3-pl_5 | 5 | 9 | 134 | 10 | 24 | 563 |
| CC_2_2_3-pl_6 | 6 | 10 | 106 | 9 | 18 | 246 |
| CC_2_2_3-pl_7 | 8 | 33837 | 214811 | |||
| CC_2_2_3-pl_8 | 8 | 247 | 2434 | 11 | 36 | 763 |
| CC_2_2_4-pl_3 | 3 | 3 | 279 | 3 | 3 | 290 |
| CC_2_2_4-pl_4 | 6 | 9 | 212 | 8 | 11 | 689 |
| CC_2_2_4-pl_5 | 9 | 77 | 2507 | 13 | 38 | 1221 |
| CC_2_2_4-pl_6 | 7 | 26 | 296 | |||
| CC_2_2_4-pl_7 | 8 | 498 | 4701 | |||
| CC_2_3_4-pl_3 | 3 | 3 | 1381 | 3 | 3 | 918 |
| CC_2_3_4-pl_4 | 6 | 10 | 1655 | 13 | 19 | 4652 |
| CC_2_3_4-pl_5 | 12 | 61 | 5848 | 13 | 24 | 5774 |
| CC_2_3_4-pl_6 | 8 | 11 | 3383 | 8 | 10 | 5843 |
| CC_2_3_4-pl_7 | 9 | 22 | 3244 | 9 | 19 | 4247 |
| CC_3_2_3-pl_3 | 3 | 3 | 196 | 3 | 3 | 330 |
| CC_3_2_3-pl_4 | 5 | 7 | 160 | 6 | 8 | 315 |
| CC_3_2_3-pl_5 | 6 | 45 | 463 | |||
| CC_3_2_3-pl_6 | 6 | 113 | 1092 | |||
| CC_3_2_3-pl_7 | 7 | 1835 | 24464 | |||
| CC_3_3_3-pl_3 | 3 | 4 | 367 | 3 | 3 | 295 |
| CC_3_3_3-pl_4 | 4 | 4 | 495 | 4 | 4 | 424 |
| CC_3_3_3-pl_5 | 7 | 22 | 534 | 8 | 11 | 670 |
| CC_3_3_3-pl_6 | ||||||
| CC_3_3_3-pl_7 | 8 | 79 | 1255 | |||
| CoinBox-pl_2 | 2 | 2 | 50 | 2 | 2 | 64 |
| CoinBox-pl_3 | 3 | 17 | 129 | 4 | 12 | 142 |
| CoinBox-pl_5 | 5 | 77 | 552 | 5 | 16 | 220 |
| CoinBox-pl_6 | 6 | 381 | 2748 | |||
| CoinBox-pl_7 | 7 | 1817 | 13110 | |||
| SC_4_1-pl_3 | 3 | 4 | 52 | 3 | 4 | 125 |
| SC_4_1-pl_5 | 5 | 17 | 99 | |||
| SC_4_2-pl_5 | 5 | 15 | 99 | |||
| SC_4_2-pl_7 | 7 | 120 | 545 | |||
| SC_4_2-pl_8 | 8 | 339 | 1316 | |||
| SC_4_3-pl_5 | 5 | 17 | 62 | |||
| SC_4_3-pl_6 | 6 | 21 | 113 | |||
| SC_4_3-pl_8 | 8 | 59 | 156 | |||
| SC_4_4-pl_5 | 5 | 17 | 87 | |||
| SC_8_10-pl_6 | 6 | 49 | 343 | 483 | 4776 | 58158 |
| SC_8_10-pl_8 | 8 | 542 | 3841 | 1586 | 19087 | 326379 |
| SC_8_10-pl_9 | 9 | 1660 | 14557 | 1586 | 19087 | 319399 |
| SC_8_10-pl_12 | 12 | 31662 | 203157 | |||
| SC_9_11-pl_4 | 4 | 8 | 141 | 6 | 13 | 172 |
| SC_9_11-pl_5 | 5 | 11 | 96 | 8 | 21 | 296 |
| SC_9_11-pl_6 | 6 | 26 | 161 | 509 | 5116 | 55844 |
| SC_9_11-pl_7 | 7 | 41 | 299 | 1286 | 15107 | 195915 |
| SC_9_11-pl_8 | 8 | 66 | 271 | 1098 | 12438 | 186999 |
| SC_9_11-pl_9 | 9 | 238 | 2158 | 1099 | 12457 | 150920 |
| SC_9_11-pl_10 | 10 | 663 | 5467 | 1100 | 12470 | 176047 |
| SC_9_11-pl_11 | 11 | 1262 | 10960 | 1100 | 12472 | 173729 |
| SC_10_8-pl_9 | ||||||
| SC_10_8-pl_10 | 10 | 10 | 97 | 10 | 10 | 112 |
| SC_10_8-pl_14 | 14 | 93 | 378 | 16 | 33 | 273 |
| SC_10_8-pl_15 | 15 | 130 | 543 | |||
| SC_10_10-pl_2 | 2 | 6 | 173 | 4 | 13 | 720 |
| SC_10_10-pl_3 | 3 | 10 | 248 | |||
| SC_10_10-pl_6 | 6 | 309 | 5998 | |||
| SC_10_10-pl_7 | 7 | 647 | 11495 | |||
| SC_10_10-pl_9 | 9 | 10 | 120 | 9 | 10 | 225 |
| SC_10_10-pl_9 | 9 | 886 | 17444 | 13 | 65 | 946 |
| SC_10_10-pl_10 | 10 | 15 | 158 | 11 | 17 | 273 |
| SC_10_10-pl_10 | 10 | 3096 | 58382 | |||
| SC_10_10-pl_13 | 13 | 15826 | 330014 | |||
| SC_10_10-pl_13 | 13 | 51 | 360 | 17 | 49 | 519 |
| SC_10_10-pl_14 | 14 | 30967 | 582175 | |||
| SC_10_10-pl_17 | 17 | 946 | 5175 | |||
| SC_10_10-pl_17 | ||||||
| () | 7 3 | 1723 6489 | 24971 88115 | 207 450 | 2322 5273 | 42635 93496 |
| - () | 6 4 | 44 232 | 971 4112 | 8 8 | 18 27 | 1290 5528 |
| () | 6 3 | 130 326 | 8372 41892 | 207 450 | 2322 5273 | 42635 93496 |
| - () | 6 4 | 17 42 | 584 2276 | 8 8 | 18 27 | 1290 5528 |
| Solved Instances | 75/79 (94.94%) | 49/79 (62.03%) | ||||
| Instance Name | HFS∗ | HFS | ||||
| [ms] | [ms] | |||||
| Assemble_B10-pl_5 | ||||||
| Assemble_B2-pl_5 | 5 | 14 | 285 | 5 | 14 | 287 |
| Assemble_B3-pl_5 | 5 | 14 | 308 | 5 | 14 | 302 |
| Assemble_B4-pl_5 | 5 | 14 | 330 | 5 | 14 | 357 |
| Assemble_B5-pl_5 | 5 | 14 | 475 | 5 | 14 | 448 |
| Assemble_B6-pl_5 | 5 | 14 | 1127 | 5 | 14 | 1190 |
| Assemble_B7-pl_5 | 5 | 14 | 9894 | 5 | 14 | 10020 |
| Assemble_B8-pl_5 | 5 | 14 | 48340 | 5 | 14 | 49503 |
| Assemble_B9-pl_5 | ||||||
| Assemble_C-pl_5 | 5 | 14 | 238 | 5 | 14 | 257 |
| CC_2_2_3-pl_3 | 3 | 9 | 181 | 4 | 10 | 184 |
| CC_2_2_3-pl_4 | 4 | 14 | 171 | 4 | 9 | 182 |
| CC_2_2_3-pl_5 | 5 | 77 | 684 | 7 | 28 | 255 |
| CC_2_2_3-pl_6 | 6 | 178 | 1532 | 7 | 25 | 291 |
| CC_2_2_3-pl_7 | 7 | 558 | 4358 | 44 | 363 | 2249 |
| CC_2_2_3-pl_8 | 8 | 2539 | 27312 | |||
| CC_2_2_4-pl_3 | 3 | 11 | 373 | 4 | 9 | 370 |
| CC_2_2_4-pl_4 | 4 | 23 | 778 | 7 | 28 | 866 |
| CC_2_2_4-pl_5 | 5 | 94 | 3483 | 12 | 70 | 2787 |
| CC_2_2_4-pl_6 | 6 | 363 | 9536 | 13 | 79 | 2764 |
| CC_2_2_4-pl_7 | 7 | 2084 | 45668 | |||
| CC_2_3_4-pl_3 | 3 | 19 | 6067 | 24 | 199 | 49682 |
| CC_2_3_4-pl_4 | 4 | 42 | 12130 | 24 | 201 | 49087 |
| CC_2_3_4-pl_5 | 5 | 127 | 45028 | 13 | 85 | 27215 |
| CC_2_3_4-pl_6 | 6 | 675 | 124726 | 29 | 256 | 56262 |
| CC_2_3_4-pl_7 | 7 | 4153 | 585278 | |||
| CC_3_2_3-pl_3 | 3 | 14 | 318 | 9 | 48 | 533 |
| CC_3_2_3-pl_4 | 4 | 47 | 689 | 5 | 16 | 254 |
| CC_3_2_3-pl_5 | 5 | 136 | 2088 | 8 | 39 | 500 |
| CC_3_2_3-pl_6 | 6 | 367 | 6443 | 6 | 21 | 323 |
| CC_3_2_3-pl_7 | 7 | 1548 | 26110 | |||
| CC_3_3_3-pl_3 | 3 | 15 | 790 | 5 | 17 | 828 |
| CC_3_3_3-pl_4 | 4 | 39 | 1337 | 4 | 10 | 599 |
| CC_3_3_3-pl_5 | 5 | 253 | 8316 | 18 | 137 | 3073 |
| CC_3_3_3-pl_6 | 6 | 1942 | 73680 | 31 | 279 | 5961 |
| CC_3_3_3-pl_7 | 7 | 11726 | 405444 | |||
| CoinBox-pl_2 | 2 | 2 | 109 | 2 | 2 | 106 |
| CoinBox-pl_3 | 3 | 17 | 228 | 4 | 12 | 201 |
| CoinBox-pl_5 | 5 | 78 | 728 | 5 | 16 | 265 |
| CoinBox-pl_6 | 6 | 381 | 3760 | |||
| CoinBox-pl_7 | 7 | 1793 | 17758 | |||
| SC_4_1-pl_3 | 3 | 4 | 96 | 3 | 4 | 109 |
| SC_4_1-pl_5 | 5 | 17 | 93 | |||
| SC_4_2-pl_5 | 5 | 30 | 184 | |||
| SC_4_2-pl_7 | 7 | 97 | 533 | |||
| SC_4_2-pl_8 | 8 | 184 | 923 | |||
| SC_4_3-pl_5 | 5 | 17 | 98 | |||
| SC_4_3-pl_6 | 6 | 21 | 93 | |||
| SC_4_3-pl_8 | 8 | 59 | 177 | |||
| SC_4_4-pl_5 | 5 | 17 | 91 | |||
| SC_8_10-pl_6 | 6 | 113 | 882 | 1951 | 24489 | 384947 |
| SC_8_10-pl_8 | 8 | 485 | 4121 | |||
| SC_8_10-pl_9 | 9 | 1772 | 16936 | |||
| SC_8_10-pl_12 | 12 | 35063 | 281914 | |||
| SC_9_11-pl_4 | 4 | 8 | 139 | 6 | 13 | 150 |
| SC_9_11-pl_5 | 5 | 11 | 122 | 8 | 21 | 189 |
| SC_9_11-pl_6 | 6 | 26 | 198 | 509 | 5116 | 48319 |
| SC_9_11-pl_7 | 7 | 41 | 326 | 1286 | 15107 | 202732 |
| SC_9_11-pl_8 | 8 | 66 | 500 | 1099 | 12447 | 152051 |
| SC_9_11-pl_9 | 9 | 237 | 2041 | |||
| SC_9_11-pl_10 | 10 | 657 | 6013 | |||
| SC_9_11-pl_11 | 11 | 1272 | 12372 | |||
| SC_10_8-pl_9 | ||||||
| SC_10_8-pl_10 | 10 | 10 | 108 | 10 | 10 | 137 |
| SC_10_8-pl_14 | 14 | 93 | 572 | 16 | 33 | 252 |
| SC_10_8-pl_15 | 15 | 130 | 822 | |||
| SC_10_10-pl_2 | 2 | 6 | 311 | 3 | 8 | 315 |
| SC_10_10-pl_3 | 3 | 10 | 347 | |||
| SC_10_10-pl_6 | 6 | 298 | 8135 | |||
| SC_10_10-pl_7 | 7 | 589 | 16064 | |||
| SC_10_10-pl_9 | 9 | 10 | 117 | 9 | 10 | 132 |
| SC_10_10-pl_9 | 9 | 1248 | 32644 | 12 | 57 | 1589 |
| SC_10_10-pl_10 | 10 | 15 | 161 | 11 | 17 | 176 |
| SC_10_10-pl_10 | 10 | 1977 | 51904 | |||
| SC_10_10-pl_13 | 13 | 17305 | 472785 | |||
| SC_10_10-pl_13 | 13 | 51 | 358 | 17 | 49 | 400 |
| SC_10_10-pl_14 | ||||||
| SC_10_10-pl_17 | 17 | 942 | 7546 | |||
| SC_10_10-pl_17 | ||||||
| () | 6 3 | 1248 4653 | 32376 101790 | 115 371 | 1292 4503 | 23015 66231 |
| - () | 6 3 | 87 445 | 2415 9514 | 7 10 | 26 63 | 808 2747 |
| () | 6 3 | 150 346 | 8724 22575 | 115 371 | 1292 4503 | 23015 66231 |
| - () | 5 2 | 32 80 | 831 3890 | 7 10 | 26 63 | 808 2747 |
| Solved Instances | 74/79 (93.67%) | 46/79 (58.23%) | ||||
F.4. Bitmask Representation Search Strategies Comparison
| Instance Name | HFS∗ | HFS | ||||
| [ms] | [ms] | |||||
| Assemble_B3-pl_5 | 5 | 14 | 164 | 5 | 14 | 378 |
| Assemble_B5-pl_5 | 5 | 14 | 223 | 5 | 14 | 323 |
| Assemble_B7-pl_5 | 5 | 14 | 4368 | 5 | 14 | 8500 |
| CC_2_2_3-pl_4 | 4 | 13 | 183 | 6 | 12 | 116 |
| CC_2_2_3-pl_6 | 6 | 62 | 424 | 7 | 17 | 228 |
| CC_2_2_4-pl_5 | 7 | 43 | 938 | 14 | 43 | 847 |
| CC_2_3_4-pl_3 | 5 | 6 | 2517 | 7 | 9 | 4534 |
| CC_2_3_4-pl_7 | 11 | 2054 | 88913 | |||
| CC_3_2_3-pl_4 | 5 | 13 | 219 | 8 | 12 | 406 |
| CC_3_2_3-pl_5 | 6 | 51 | 888 | 21 | 3442 | 92075 |
| CC_3_3_3-pl_4 | 6 | 19 | 1730 | 7 | 8 | 788 |
| CoinBox-pl_3 | 3 | 17 | 220 | 4 | 12 | 160 |
| CoinBox-pl_6 | 6 | 381 | 3534 | |||
| SC_4_1-pl_3 | 3 | 4 | 51 | 3 | 4 | 57 |
| SC_4_2-pl_7 | 7 | 120 | 609 | |||
| SC_4_3-pl_5 | 5 | 17 | 116 | |||
| SC_8_10-pl_6 | 6 | 49 | 315 | 483 | 4776 | 50264 |
| SC_9_11-pl_6 | 6 | 26 | 179 | |||
| SC_9_11-pl_8 | 8 | 66 | 388 | |||
| SC_10_8-pl_10 | 10 | 10 | 174 | 10 | 10 | 153 |
| SC_10_8-pl_15 | 15 | 130 | 824 | |||
| SC_10_10-pl_6 | 6 | 309 | 8676 | |||
| SC_10_10-pl_7 | 7 | 647 | 17999 | |||
| SC_10_10-pl_10 | 10 | 3096 | 80030 | |||
| SC_10_10-pl_13 | 13 | 51 | 431 | 17 | 49 | 406 |
| () | 7 3 | 289 705 | 8565 22733 | 40 118 | 562 1412 | 10616 25045 |
| - () | 6 2 | 40 106 | 748 2298 | 7 7 | 14 19 | 482 2496 |
| () | 6 2 | 25 19 | 856 1148 | 40 118 | 562 1412 | 10616 25045 |
| - () | 5 1 | 18 33 | 389 712 | 7 7 | 14 19 | 482 2496 |
| Solved Instances | 25/25 (100.00%) | 15/25 (60.00%) | ||||
| Instance Name | HFS∗ | HFS | ||||
| [ms] | [ms] | |||||
| Assemble_B10-pl_5 | ||||||
| Assemble_B2-pl_5 | 5 | 14 | 264 | 5 | 14 | 269 |
| Assemble_B4-pl_5 | 5 | 14 | 275 | 5 | 14 | 376 |
| Assemble_B6-pl_5 | 5 | 14 | 804 | 5 | 14 | 1025 |
| Assemble_B8-pl_5 | 5 | 14 | 24655 | 5 | 14 | 35914 |
| Assemble_B9-pl_5 | 5 | 14 | 295842 | 5 | 14 | 465844 |
| Assemble_C-pl_5 | 5 | 14 | 124 | 5 | 14 | 174 |
| CC_2_2_3-pl_3 | 3 | 9 | 111 | 4 | 7 | 103 |
| CC_2_2_3-pl_5 | 5 | 17 | 272 | 5 | 14 | 156 |
| CC_2_2_3-pl_7 | 7 | 163 | 2390 | 11 | 2150 | 6290 |
| CC_2_2_3-pl_8 | 8 | 1327 | 18235 | |||
| CC_2_2_4-pl_3 | 3 | 3 | 217 | 3 | 3 | 202 |
| CC_2_2_4-pl_4 | 6 | 7 | 324 | 6 | 6 | 390 |
| CC_2_2_4-pl_6 | 6 | 21 | 1251 | 55 | 1062 | 10165 |
| CC_2_2_4-pl_7 | 8 | 129 | 3989 | |||
| CC_2_3_4-pl_4 | 5 | 5 | 1995 | 5 | 5 | 1126 |
| CC_2_3_4-pl_5 | 10 | 132 | 20181 | |||
| CC_2_3_4-pl_6 | 11 | 1269 | 65775 | |||
| CC_3_2_3-pl_3 | 3 | 4 | 228 | 3 | 3 | 236 |
| CC_3_2_3-pl_6 | 6 | 74 | 1201 | 19 | 311 | 12446 |
| CC_3_2_3-pl_7 | 8 | 185 | 3058 | |||
| CC_3_3_3-pl_3 | 3 | 3 | 218 | 3 | 3 | 212 |
| CC_3_3_3-pl_5 | 6 | 26 | 1975 | 8 | 12 | 2513 |
| CC_3_3_3-pl_6 | 8 | 226 | 13395 | 10 | 18 | 1864 |
| CC_3_3_3-pl_7 | 8 | 1476 | 59481 | |||
| CoinBox-pl_2 | 2 | 2 | 52 | 2 | 2 | 71 |
| CoinBox-pl_5 | 5 | 77 | 425 | 5 | 16 | 156 |
| CoinBox-pl_7 | 7 | 1817 | 10395 | |||
| SC_4_1-pl_5 | 5 | 17 | 125 | |||
| SC_4_2-pl_5 | 5 | 15 | 132 | |||
| SC_4_2-pl_8 | 8 | 339 | 1639 | |||
| SC_4_3-pl_6 | 6 | 21 | 162 | |||
| SC_4_3-pl_8 | 8 | 59 | 186 | |||
| SC_4_4-pl_5 | 5 | 17 | 125 | |||
| SC_8_10-pl_8 | 8 | 542 | 5157 | 1586 | 19087 | 329262 |
| SC_8_10-pl_9 | 9 | 1660 | 17268 | 1586 | 19087 | 321429 |
| SC_8_10-pl_12 | 12 | 31662 | 199243 | |||
| SC_9_11-pl_4 | 4 | 8 | 143 | 6 | 13 | 192 |
| SC_9_11-pl_5 | 5 | 11 | 152 | 8 | 21 | 315 |
| SC_9_11-pl_7 | 7 | 41 | 431 | |||
| SC_9_11-pl_9 | 9 | 238 | 1895 | |||
| SC_9_11-pl_10 | 10 | 663 | 6151 | |||
| SC_9_11-pl_11 | 11 | 1262 | 15073 | |||
| SC_10_8-pl_9 | ||||||
| SC_10_8-pl_14 | 14 | 93 | 481 | 16 | 33 | 390 |
| SC_10_10-pl_2 | 2 | 6 | 280 | 4 | 13 | 386 |
| SC_10_10-pl_3 | 3 | 10 | 239 | |||
| SC_10_10-pl_9 | 9 | 10 | 188 | 9 | 10 | 114 |
| SC_10_10-pl_9 | 9 | 886 | 20614 | 13 | 65 | 1442 |
| SC_10_10-pl_10 | 10 | 15 | 191 | 11 | 17 | 220 |
| SC_10_10-pl_13 | 13 | 15826 | 285398 | |||
| SC_10_10-pl_14 | ||||||
| SC_10_10-pl_17 | 17 | 946 | 7240 | |||
| SC_10_10-pl_17 | ||||||
| () | 7 3 | 1228 4886 | 21793 62537 | 118 400 | 1450 4819 | 41148 114657 |
| - () | 6 4 | 66 300 | 1758 9389 | 7 6 | 15 11 | 731 2311 |
| () | 6 3 | 136 342 | 13465 53762 | 118 400 | 1450 4819 | 41148 114657 |
| - () | 5 2 | 18 66 | 681 1778 | 7 6 | 15 11 | 731 2311 |
| Solved Instances | 50/54 (92.59%) | 29/54 (53.70%) | ||||
| Instance Name | HFS∗ | HFS | ||||
| [ms] | [ms] | |||||
| Assemble_B10-pl_5 | ||||||
| Assemble_B2-pl_5 | 5 | 14 | 264 | 5 | 14 | 269 |
| Assemble_B3-pl_5 | 5 | 14 | 164 | 5 | 14 | 378 |
| Assemble_B4-pl_5 | 5 | 14 | 275 | 5 | 14 | 376 |
| Assemble_B5-pl_5 | 5 | 14 | 223 | 5 | 14 | 323 |
| Assemble_B6-pl_5 | 5 | 14 | 804 | 5 | 14 | 1025 |
| Assemble_B7-pl_5 | 5 | 14 | 4368 | 5 | 14 | 8500 |
| Assemble_B8-pl_5 | 5 | 14 | 24655 | 5 | 14 | 35914 |
| Assemble_B9-pl_5 | 5 | 14 | 295842 | 5 | 14 | 465844 |
| Assemble_C-pl_5 | 5 | 14 | 124 | 5 | 14 | 174 |
| CC_2_2_3-pl_3 | 3 | 9 | 111 | 4 | 7 | 103 |
| CC_2_2_3-pl_4 | 4 | 13 | 183 | 6 | 12 | 116 |
| CC_2_2_3-pl_5 | 5 | 17 | 272 | 5 | 14 | 156 |
| CC_2_2_3-pl_6 | 6 | 62 | 424 | 7 | 17 | 228 |
| CC_2_2_3-pl_7 | 7 | 163 | 2390 | 11 | 2150 | 6290 |
| CC_2_2_3-pl_8 | 8 | 1327 | 18235 | |||
| CC_2_2_4-pl_3 | 3 | 3 | 217 | 3 | 3 | 202 |
| CC_2_2_4-pl_4 | 6 | 7 | 324 | 6 | 6 | 390 |
| CC_2_2_4-pl_5 | 7 | 43 | 938 | 14 | 43 | 847 |
| CC_2_2_4-pl_6 | 6 | 21 | 1251 | 55 | 1062 | 10165 |
| CC_2_2_4-pl_7 | 8 | 129 | 3989 | |||
| CC_2_3_4-pl_3 | 5 | 6 | 2517 | 7 | 9 | 4534 |
| CC_2_3_4-pl_4 | 5 | 5 | 1995 | 5 | 5 | 1126 |
| CC_2_3_4-pl_5 | 10 | 132 | 20181 | |||
| CC_2_3_4-pl_6 | 11 | 1269 | 65775 | |||
| CC_2_3_4-pl_7 | 11 | 2054 | 88913 | |||
| CC_3_2_3-pl_3 | 3 | 4 | 228 | 3 | 3 | 236 |
| CC_3_2_3-pl_4 | 5 | 13 | 219 | 8 | 12 | 406 |
| CC_3_2_3-pl_5 | 6 | 51 | 888 | 21 | 3442 | 92075 |
| CC_3_2_3-pl_6 | 6 | 74 | 1201 | 19 | 311 | 12446 |
| CC_3_2_3-pl_7 | 8 | 185 | 3058 | |||
| CC_3_3_3-pl_3 | 3 | 3 | 218 | 3 | 3 | 212 |
| CC_3_3_3-pl_4 | 6 | 19 | 1730 | 7 | 8 | 788 |
| CC_3_3_3-pl_5 | 6 | 26 | 1975 | 8 | 12 | 2513 |
| CC_3_3_3-pl_6 | 8 | 226 | 13395 | 10 | 18 | 1864 |
| CC_3_3_3-pl_7 | 8 | 1476 | 59481 | |||
| CoinBox-pl_2 | 2 | 2 | 52 | 2 | 2 | 71 |
| CoinBox-pl_3 | 3 | 17 | 220 | 4 | 12 | 160 |
| CoinBox-pl_5 | 5 | 77 | 425 | 5 | 16 | 156 |
| CoinBox-pl_6 | 6 | 381 | 3534 | |||
| CoinBox-pl_7 | 7 | 1817 | 10395 | |||
| SC_4_1-pl_3 | 3 | 4 | 51 | 3 | 4 | 57 |
| SC_4_1-pl_5 | 5 | 17 | 125 | |||
| SC_4_2-pl_5 | 5 | 15 | 132 | |||
| SC_4_2-pl_7 | 7 | 120 | 609 | |||
| SC_4_2-pl_8 | 8 | 339 | 1639 | |||
| SC_4_3-pl_5 | 5 | 17 | 116 | |||
| SC_4_3-pl_6 | 6 | 21 | 162 | |||
| SC_4_3-pl_8 | 8 | 59 | 186 | |||
| SC_4_4-pl_5 | 5 | 17 | 125 | |||
| SC_8_10-pl_6 | 6 | 49 | 315 | 483 | 4776 | 50264 |
| SC_8_10-pl_8 | 8 | 542 | 5157 | 1586 | 19087 | 329262 |
| SC_8_10-pl_9 | 9 | 1660 | 17268 | 1586 | 19087 | 321429 |
| SC_8_10-pl_12 | 12 | 31662 | 199243 | |||
| SC_9_11-pl_4 | 4 | 8 | 143 | 6 | 13 | 192 |
| SC_9_11-pl_5 | 5 | 11 | 152 | 8 | 21 | 315 |
| SC_9_11-pl_6 | 6 | 26 | 179 | |||
| SC_9_11-pl_7 | 7 | 41 | 431 | |||
| SC_9_11-pl_8 | 8 | 66 | 388 | |||
| SC_9_11-pl_9 | 9 | 238 | 1895 | |||
| SC_9_11-pl_10 | 10 | 663 | 6151 | |||
| SC_9_11-pl_11 | 11 | 1262 | 15073 | |||
| SC_10_8-pl_9 | ||||||
| SC_10_8-pl_10 | 10 | 10 | 174 | 10 | 10 | 153 |
| SC_10_8-pl_14 | 14 | 93 | 481 | 16 | 33 | 390 |
| SC_10_8-pl_15 | 15 | 130 | 824 | |||
| SC_10_10-pl_2 | 2 | 6 | 280 | 4 | 13 | 386 |
| SC_10_10-pl_3 | 3 | 10 | 239 | |||
| SC_10_10-pl_6 | 6 | 309 | 8676 | |||
| SC_10_10-pl_7 | 7 | 647 | 17999 | |||
| SC_10_10-pl_9 | 9 | 10 | 188 | 9 | 10 | 114 |
| SC_10_10-pl_9 | 9 | 886 | 20614 | 13 | 65 | 1442 |
| SC_10_10-pl_10 | 10 | 15 | 191 | 11 | 17 | 220 |
| SC_10_10-pl_10 | 10 | 3096 | 80030 | |||
| SC_10_10-pl_13 | 13 | 15826 | 285398 | |||
| SC_10_10-pl_13 | 13 | 51 | 431 | 17 | 49 | 406 |
| SC_10_10-pl_14 | ||||||
| SC_10_10-pl_17 | 17 | 946 | 7240 | |||
| SC_10_10-pl_17 | ||||||
| () | 7 3 | 915 4034 | 17383 53089 | 91 334 | 1147 4021 | 30739 95330 |
| - () | 6 3 | 54 218 | 1255 5436 | 7 6 | 14 14 | 652 2819 |
| () | 6 3 | 98 283 | 9167 44059 | 91 334 | 1147 4021 | 30739 95330 |
| - () | 5 1 | 18 40 | 529 1581 | 7 6 | 14 14 | 652 2819 |
| Solved Instances | 75/79 (94.94%) | 44/79 (55.70%) | ||||
Appendix G Experimental Results
In what follows, we present the results for all the experiments we conducted.
We will use the following abbreviations:
GNN: denotes our primary contribution, deep equipped with A∗ search and GNN-based heuristics.
BFS: deep using breadth-first search.
-EFP: a planner that also incorporates heuristics, albeit not derived through learning techniques (DBLP:conf/prima/FabianoPSP24). Following DBLP:conf/prima/FabianoPSP24, we will use the following to distinguish between the various heuristics used by -EFP
SUB: Heuristics that associates a higher evaluation to e-states that satisfy more sub-goals.
C_PG: Heuristics that emulates the classical Planning Graph by deriving the “importance” of each belief formula (its distance from the goal level) and then each e-state is characterized by the sum of the derived belief formulae scores. In particular C_PG reflects the hAdd heuristics in MEP.
L_PG: Heuristics that calculates the score of an e-state by constructing a Planning Graph from it (as initial state) and calculating its length—the shorter the better—of the constructed e-PG. This behavior is similar to the one adopted by the heuristics hFF in classical planning.
S_PG: This heuristics is simply an execution of C_PG on every e-state.
: the number of nodes expanded during search, reflecting the informativeness of the heuristics.
: the length of the plan found.
: the solving time (in milliseconds).
: Interquartile Mean, used as an aggregate performance metric.
-: Interquartile Range, reported as a measure of variability.
: arithmetic mean, reported as a baseline aggregate metric.
: standard deviation, used to quantify variability in the data.
: indicates that the aggregate value is computed over all instances solved by that approach.
: indicates that the aggregate value is computed only over instances solved by all approaches in the comparison.
: indicates a timeout (after 600 seconds).
: indicates a missing value due to the problem not being solved within the timeout.
Since each set of experiments includes both “Train” and “Test” problems, we report results for each in separate tables. This distinction is indicated in the table captions using either Train or Test.
G.1. Experimental Setup #1 – Standard Benchmarks
| Instance Name | GNN | BFS | ||||
| [ms] | [ms] | |||||
| Assemble_B3-pl_5 | 5 | 10 | 151 | 5 | 14 | 39 |
| Assemble_B5-pl_5 | 5 | 10 | 124 | 5 | 14 | 55 |
| Assemble_B7-pl_5 | 5 | 10 | 891 | 5 | 14 | 1057 |
| () | 5 0 | 10 0 | 389 355 | 5 0 | 14 0 | 384 476 |
| - () | 5 0 | 10 0 | 151 384 | 5 0 | 14 0 | 55 509 |
| () | 5 0 | 10 0 | 389 355 | 5 0 | 14 0 | 384 476 |
| - () | 5 0 | 10 0 | 151 384 | 5 0 | 14 0 | 55 509 |
| Solved Instances | 3/3 (100.00%) | 3/3 (100.00%) | ||||
| Instance Name | GNN | BFS | ||||
| [ms] | [ms] | |||||
| Assemble_B2-pl_5 | 5 | 10 | 364 | 5 | 14 | 37 |
| Assemble_B4-pl_5 | 5 | 10 | 122 | 5 | 14 | 40 |
| Assemble_B6-pl_5 | 5 | 10 | 185 | 5 | 14 | 123 |
| Assemble_B8-pl_5 | 5 | 10 | 4270 | 5 | 14 | 5145 |
| Assemble_B9-pl_5 | 5 | 10 | 49596 | 5 | 14 | 65910 |
| Assemble_B10-pl_5 | ||||||
| Assemble_C-pl_5 | 5 | 10 | 140 | 5 | 14 | 38 |
| () | 5 0 | 10 0 | 9113 18166 | 5 0 | 14 0 | 11882 24233 |
| - () | 5 0 | 10 0 | 274 3142 | 5 0 | 14 0 | 82 3851 |
| () | 5 0 | 10 0 | 9113 18166 | 5 0 | 14 0 | 11882 24233 |
| - () | 5 0 | 10 0 | 274 3142 | 5 0 | 14 0 | 82 3851 |
| Solved Instances | 6/7 (85.71%) | 6/7 (85.71%) | ||||
| Instance Name | GNN | BFS | ||||
| [ms] | [ms] | |||||
| CC_2_2_3-pl_4 | 5 | 8 | 120 | 4 | 17 | 20 |
| CC_2_2_3-pl_6 | 8 | 67 | 284 | 6 | 287 | 300 |
| CC_2_2_4-pl_5 | 7 | 22 | 217 | 5 | 170 | 632 |
| CC_2_3_4-pl_3 | 3 | 19 | 2640 | 3 | 6 | 253 |
| CC_2_3_4-pl_7 | 7 | 8907 | 93302 | |||
| CC_3_2_3-pl_4 | 5 | 7 | 99 | 4 | 30 | 60 |
| CC_3_2_3-pl_5 | 8 | 23 | 179 | 5 | 178 | 317 |
| CC_3_3_3-pl_4 | 7 | 17 | 634 | 4 | 39 | 120 |
| () | 6 2 | 23 19 | 596 851 | 5 1 | 1204 2913 | 11876 30777 |
| - () | 6 2 | 19 10 | 227 310 | 4 1 | 104 178 | 248 291 |
| () | 6 2 | 23 19 | 596 851 | 4 1 | 104 100 | 243 192 |
| - () | 6 2 | 19 10 | 227 310 | 4 1 | 80 150 | 224 218 |
| Solved Instances | 7/8 (87.50%) | 8/8 (100.00%) | ||||
| Instance Name | GNN | BFS | ||||
| [ms] | [ms] | |||||
| CC_2_2_3-pl_3 | 4 | 7 | 155 | 3 | 9 | 13 |
| CC_2_2_3-pl_5 | 8 | 226 | 1030 | 5 | 78 | 92 |
| CC_2_2_3-pl_7 | 9 | 271 | 972 | 7 | 1065 | 1111 |
| CC_2_2_3-pl_8 | 8 | 2373 | 7418 | 8 | 2065 | 2126 |
| CC_2_2_4-pl_3 | 3 | 11 | 300 | 3 | 6 | 42 |
| CC_2_2_4-pl_4 | 5 | 10 | 209 | 4 | 24 | 115 |
| CC_2_2_4-pl_6 | 7 | 10 | 185 | 6 | 822 | 1505 |
| CC_2_2_4-pl_7 | 8 | 280 | 1857 | 7 | 4154 | 7004 |
| CC_2_3_4-pl_4 | 4 | 16 | 2064 | 4 | 29 | 1173 |
| CC_2_3_4-pl_5 | 13 | 53 | 2250 | 5 | 275 | 7629 |
| CC_2_3_4-pl_6 | 9 | 20 | 1213 | 6 | 1601 | 22473 |
| CC_3_2_3-pl_3 | 3 | 3 | 104 | 3 | 8 | 20 |
| CC_3_2_3-pl_6 | 6 | 8 | 127 | 6 | 469 | 830 |
| CC_3_2_3-pl_7 | 8 | 42 | 325 | 7 | 3345 | 7256 |
| CC_3_3_3-pl_3 | 3 | 3 | 194 | 3 | 8 | 68 |
| CC_3_3_3-pl_5 | 10 | 745 | 14505 | 5 | 538 | 1347 |
| CC_3_3_3-pl_6 | 8 | 200 | 2970 | 6 | 1643 | 4670 |
| CC_3_3_3-pl_7 | 9 | 456 | 5932 | 7 | 12149 | 30566 |
| () | 7 3 | 263 548 | 2323 3566 | 5 2 | 1572 2824 | 4891 8173 |
| - () | 7 4 | 65 250 | 996 2006 | 5 3 | 610 1607 | 1610 6323 |
| () | 7 3 | 263 548 | 2323 3566 | 5 2 | 1572 2824 | 4891 8173 |
| - () | 7 4 | 65 250 | 996 2006 | 5 3 | 610 1607 | 1610 6323 |
| Solved Instances | 18/18 (100.00%) | 18/18 (100.00%) | ||||
| Instance Name | GNN | BFS | ||||
| [ms] | [ms] | |||||
| CoinBox-pl_3 | 3 | 17 | 139 | 3 | 11 | 22 |
| CoinBox-pl_6 | 6 | 359 | 1425 | 6 | 912 | 1057 |
| () | 4 2 | 188 171 | 782 643 | 4 2 | 462 450 | 540 518 |
| - () | 4 2 | 188 171 | 782 643 | 4 2 | 462 450 | 540 518 |
| () | 4 2 | 188 171 | 782 643 | 4 2 | 462 450 | 540 518 |
| - () | 4 2 | 188 171 | 782 643 | 4 2 | 462 450 | 540 518 |
| Solved Instances | 2/2 (100.00%) | 2/2 (100.00%) | ||||
| Instance Name | GNN | BFS | ||||
| [ms] | [ms] | |||||
| CoinBox-pl_2 | 2 | 2 | 84 | 2 | 2 | 8 |
| CoinBox-pl_5 | 5 | 75 | 279 | 5 | 102 | 118 |
| CoinBox-pl_7 | 7 | 1721 | 7102 | 7 | 2523 | 3214 |
| () | 5 2 | 599 794 | 2488 3263 | 5 2 | 876 1166 | 1113 1486 |
| - () | 5 2 | 75 860 | 279 3509 | 5 2 | 102 1260 | 118 1603 |
| () | 5 2 | 599 794 | 2488 3263 | 5 2 | 876 1166 | 1113 1486 |
| - () | 5 2 | 75 860 | 279 3509 | 5 2 | 102 1260 | 118 1603 |
| Solved Instances | 3/3 (100.00%) | 3/3 (100.00%) | ||||
| Instance Name | GNN | BFS | ||||
| [ms] | [ms] | |||||
| Grapevine_3-pl_3 | 3 | 18 | 162 | 3 | 31 | 117 |
| Grapevine_3-pl_4 | 6 | 38 | 200 | 4 | 160 | 667 |
| Grapevine_4-pl_2 | 2 | 3 | 131 | 2 | 7 | 101 |
| Grapevine_5-pl_4 | 7 | 68 | 58927 | 4 | 616 | 61732 |
| () | 4 2 | 32 24 | 14855 25445 | 3 1 | 204 245 | 15654 26604 |
| - () | 4 4 | 28 31 | 181 14728 | 4 1 | 96 249 | 392 15820 |
| () | 4 2 | 32 24 | 14855 25445 | 3 1 | 204 245 | 15654 26604 |
| - () | 4 4 | 28 31 | 181 14728 | 4 1 | 96 249 | 392 15820 |
| Solved Instances | 4/4 (100.00%) | 4/4 (100.00%) | ||||
| Instance Name | GNN | BFS | ||||
| [ms] | [ms] | |||||
| Grapevine_3-pl_2 | 2 | 4 | 114 | 2 | 6 | 22 |
| Grapevine_3-pl_5 | 5 | 823 | 2028 | |||
| Grapevine_3-pl_6 | 6 | 2118 | 4353 | |||
| Grapevine_3-pl_7 | 9 | 10414 | 144193 | 7 | 12008 | 19763 |
| Grapevine_4-pl_3 | 6 | 332 | 5342 | 3 | 40 | 612 |
| Grapevine_4-pl_4 | 7 | 190 | 10743 | 4 | 233 | 1429 |
| Grapevine_4-pl_5 | 6 | 90 | 6738 | 5 | 1449 | 7077 |
| Grapevine_4-pl_6 | 7 | 581 | 33904 | 6 | 4055 | 18234 |
| Grapevine_5-pl_2 | 3 | 3 | 398 | 2 | 8 | 815 |
| Grapevine_5-pl_3 | 6 | 17 | 3224 | 3 | 71 | 11056 |
| Grapevine_5-pl_5 | ||||||
| Grapevine_5-pl_6 | ||||||
| () | 6 2 | 1454 3392 | 25582 45969 | 4 2 | 2081 3531 | 6539 7028 |
| - () | 6 2 | 157 380 | 6512 14016 | 4 3 | 644 1903 | 3722 9093 |
| () | 6 2 | 1454 3392 | 25582 45969 | 4 2 | 2234 3919 | 7376 7609 |
| - () | 6 2 | 157 380 | 6512 14016 | 4 2 | 448 2068 | 5094 12086 |
| Solved Instances | 8/12 (66.67%) | 10/12 (83.33%) | ||||
| Instance Name | GNN | BFS | ||||
| [ms] | [ms] | |||||
| SC_4_1-pl_3 | 3 | 3 | 56 | 3 | 4 | 2 |
| SC_4_2-pl_7 | 7 | 729 | 549 | 7 | 136 | 106 |
| SC_4_3-pl_5 | 5 | 31 | 81 | 5 | 11 | 4 |
| SC_8_10-pl_6 | 10 | 24 | 131 | 6 | 113 | 148 |
| SC_9_11-pl_6 | 6 | 19 | 99 | 6 | 17 | 20 |
| SC_9_11-pl_8 | 10 | 15 | 96 | 8 | 65 | 91 |
| SC_10_8-pl_10 | 10 | 10 | 60 | 10 | 10 | 8 |
| SC_10_8-pl_15 | 17 | 23 | 88 | 15 | 96 | 92 |
| SC_10_10-pl_13 | 14 | 33 | 165 | 13 | 52 | 49 |
| () | 9 4 | 99 223 | 147 146 | 8 4 | 56 47 | 58 50 |
| - () | 9 4 | 22 16 | 99 50 | 7 4 | 48 85 | 52 84 |
| () | 9 4 | 99 223 | 147 146 | 8 4 | 56 47 | 58 50 |
| - () | 9 4 | 22 16 | 99 50 | 7 4 | 48 85 | 52 84 |
| Solved Instances | 9/9 (100.00%) | 9/9 (100.00%) | ||||
| Instance Name | GNN | BFS | ||||
| [ms] | [ms] | |||||
| SC_4_1-pl_5 | 5 | 26 | 100 | 5 | 11 | 4 |
| SC_4_2-pl_5 | 5 | 549 | 431 | 5 | 22 | 24 |
| SC_4_2-pl_8 | 11 | 1116 | 942 | 8 | 318 | 258 |
| SC_4_3-pl_6 | 6 | 26 | 77 | 6 | 21 | 6 |
| SC_4_3-pl_8 | 8 | 75 | 112 | 8 | 66 | 16 |
| SC_4_4-pl_5 | 5 | 13 | 64 | 5 | 11 | 4 |
| SC_8_10-pl_8 | 12 | 802 | 3550 | 8 | 413 | 571 |
| SC_8_10-pl_9 | 12 | 1175 | 5247 | 9 | 1399 | 2267 |
| SC_8_10-pl_12 | 13 | 12336 | 42301 | 12 | 22486 | 34954 |
| SC_9_11-pl_4 | 4 | 4 | 65 | 4 | 6 | 8 |
| SC_9_11-pl_5 | 6 | 6 | 64 | 5 | 10 | 12 |
| SC_9_11-pl_7 | 7 | 9 | 75 | 7 | 32 | 30 |
| SC_9_11-pl_9 | 11 | 23 | 141 | 9 | 146 | 216 |
| SC_9_11-pl_10 | 12 | 30 | 199 | 10 | 355 | 590 |
| SC_9_11-pl_11 | 12 | 40 | 197 | 11 | 907 | 1617 |
| SC_10_8-pl_9 | ||||||
| SC_10_8-pl_14 | 16 | 19 | 83 | 14 | 51 | 40 |
| SC_10_10-pl_9 | 9 | 9 | 65 | 9 | 10 | 7 |
| SC_10_10-pl_10 | 11 | 11 | 72 | 10 | 12 | 9 |
| SC_10_10-pl_17 | 18 | 189 | 738 | 17 | 1075 | 1567 |
| () | 10 4 | 866 2729 | 2870 9389 | 9 3 | 1440 4977 | 2221 7743 |
| - () | 10 6 | 49 357 | 157 511 | 8 4 | 114 372 | 131 572 |
| () | 10 4 | 866 2729 | 2870 9389 | 9 3 | 1440 4977 | 2221 7743 |
| - () | 10 6 | 49 357 | 157 511 | 8 4 | 114 372 | 131 572 |
| Solved Instances | 19/20 (95.00%) | 19/20 (95.00%) | ||||
| Instance Name | GNN | BFS | ||||
| [ms] | [ms] | |||||
| SC_R_10_10-pl_6 | 6 | 298 | 1987 | 6 | 333 | 1635 |
| SC_R_10_10-pl_7 | 7 | 589 | 3870 | 7 | 792 | 3951 |
| SC_R_10_10-pl_10 | 10 | 1977 | 11786 | 10 | 2431 | 11831 |
| () | 8 2 | 955 733 | 5881 4246 | 8 2 | 1185 901 | 5806 4364 |
| - () | 7 2 | 589 840 | 3870 4900 | 7 2 | 792 1049 | 3951 5098 |
| () | 8 2 | 955 733 | 5881 4246 | 8 2 | 1185 901 | 5806 4364 |
| - () | 7 2 | 589 840 | 3870 4900 | 7 2 | 792 1049 | 3951 5098 |
| Solved Instances | 3/3 (100.00%) | 3/3 (100.00%) | ||||
| Instance Name | GNN | BFS | ||||
| [ms] | [ms] | |||||
| SC_R_10_10-pl_2 | 2 | 6 | 125 | 2 | 5 | 45 |
| SC_R_10_10-pl_3 | 3 | 10 | 109 | 3 | 17 | 88 |
| SC_R_10_10-pl_9 | 9 | 1248 | 7447 | 9 | 1112 | 6079 |
| SC_R_10_10-pl_13 | 13 | 17305 | 106761 | 13 | 22625 | 120270 |
| SC_R_10_10-pl_14 | 14 | 38714 | 239252 | 14 | 47255 | 239197 |
| SC_R_10_10-pl_17 | ||||||
| () | 8 5 | 11457 15123 | 70739 93461 | 8 5 | 14203 18642 | 73136 94840 |
| - () | 8 10 | 6188 17295 | 38111 106636 | 8 10 | 7918 22608 | 42146 120182 |
| () | 8 5 | 11457 15123 | 70739 93461 | 8 5 | 14203 18642 | 73136 94840 |
| - () | 8 10 | 6188 17295 | 38111 106636 | 8 10 | 7918 22608 | 42146 120182 |
| Solved Instances | 5/6 (83.33%) | 5/6 (83.33%) | ||||
| Instance Name | GNN | BFS | ||||
| [ms] | [ms] | |||||
| Assemble_B3-pl_5 | 5 | 10 | 151 | 5 | 14 | 39 |
| Assemble_B5-pl_5 | 5 | 10 | 124 | 5 | 14 | 55 |
| Assemble_B7-pl_5 | 5 | 10 | 891 | 5 | 14 | 1057 |
| CC_2_2_3-pl_4 | 5 | 8 | 120 | 4 | 17 | 20 |
| CC_2_2_3-pl_6 | 8 | 67 | 284 | 6 | 287 | 300 |
| CC_2_2_4-pl_5 | 7 | 22 | 217 | 5 | 170 | 632 |
| CC_2_3_4-pl_3 | 3 | 19 | 2640 | 3 | 6 | 253 |
| CC_2_3_4-pl_7 | 7 | 8907 | 93302 | |||
| CC_3_2_3-pl_4 | 5 | 7 | 99 | 4 | 30 | 60 |
| CC_3_2_3-pl_5 | 8 | 23 | 179 | 5 | 178 | 317 |
| CC_3_3_3-pl_4 | 7 | 17 | 634 | 4 | 39 | 120 |
| CoinBox-pl_3 | 3 | 17 | 139 | 3 | 11 | 22 |
| CoinBox-pl_6 | 6 | 359 | 1425 | 6 | 912 | 1057 |
| Grapevine_3-pl_3 | 3 | 18 | 162 | 3 | 31 | 117 |
| Grapevine_3-pl_4 | 6 | 38 | 200 | 4 | 160 | 667 |
| Grapevine_4-pl_2 | 2 | 3 | 131 | 2 | 7 | 101 |
| Grapevine_5-pl_4 | 7 | 68 | 58927 | 4 | 616 | 61732 |
| SC_R_10_10-pl_6 | 6 | 298 | 1987 | 6 | 333 | 1635 |
| SC_R_10_10-pl_7 | 7 | 589 | 3870 | 7 | 792 | 3951 |
| SC_R_10_10-pl_10 | 10 | 1977 | 11786 | 10 | 2431 | 11831 |
| SC_4_1-pl_3 | 3 | 3 | 56 | 3 | 4 | 2 |
| SC_4_2-pl_7 | 7 | 729 | 549 | 7 | 136 | 106 |
| SC_4_3-pl_5 | 5 | 31 | 81 | 5 | 11 | 4 |
| SC_8_10-pl_6 | 10 | 24 | 131 | 6 | 113 | 148 |
| SC_9_11-pl_6 | 6 | 19 | 99 | 6 | 17 | 20 |
| SC_9_11-pl_8 | 10 | 15 | 96 | 8 | 65 | 91 |
| SC_10_8-pl_10 | 10 | 10 | 60 | 10 | 10 | 8 |
| SC_10_8-pl_15 | 17 | 23 | 88 | 15 | 96 | 92 |
| SC_10_10-pl_13 | 14 | 33 | 165 | 13 | 52 | 49 |
| () | 7 3 | 159 393 | 3046 10989 | 6 3 | 534 1652 | 6131 19990 |
| - () | 6 3 | 20 35 | 228 584 | 5 3 | 72 164 | 207 618 |
| () | 7 3 | 159 393 | 3046 10989 | 6 3 | 234 483 | 3017 11523 |
| - () | 6 3 | 20 35 | 228 584 | 5 2 | 65 158 | 174 594 |
| Solved Instances | 28/29 (96.55%) | 29/29 (100.00%) | ||||
| Instance Name | GNN | BFS | ||||
| [ms] | [ms] | |||||
| Assemble_B10-pl_5 | ||||||
| Assemble_B2-pl_5 | 5 | 10 | 364 | 5 | 14 | 37 |
| Assemble_B4-pl_5 | 5 | 10 | 122 | 5 | 14 | 40 |
| Assemble_B6-pl_5 | 5 | 10 | 185 | 5 | 14 | 123 |
| Assemble_B8-pl_5 | 5 | 10 | 4270 | 5 | 14 | 5145 |
| Assemble_B9-pl_5 | 5 | 10 | 49596 | 5 | 14 | 65910 |
| Assemble_C-pl_5 | 5 | 10 | 140 | 5 | 14 | 38 |
| CC_2_2_3-pl_3 | 4 | 7 | 155 | 3 | 9 | 13 |
| CC_2_2_3-pl_5 | 8 | 226 | 1030 | 5 | 78 | 92 |
| CC_2_2_3-pl_7 | 9 | 271 | 972 | 7 | 1065 | 1111 |
| CC_2_2_3-pl_8 | 8 | 2373 | 7418 | 8 | 2065 | 2126 |
| CC_2_2_4-pl_3 | 3 | 11 | 300 | 3 | 6 | 42 |
| CC_2_2_4-pl_4 | 5 | 10 | 209 | 4 | 24 | 115 |
| CC_2_2_4-pl_6 | 7 | 10 | 185 | 6 | 822 | 1505 |
| CC_2_2_4-pl_7 | 8 | 280 | 1857 | 7 | 4154 | 7004 |
| CC_2_3_4-pl_4 | 4 | 16 | 2064 | 4 | 29 | 1173 |
| CC_2_3_4-pl_5 | 13 | 53 | 2250 | 5 | 275 | 7629 |
| CC_2_3_4-pl_6 | 9 | 20 | 1213 | 6 | 1601 | 22473 |
| CC_3_2_3-pl_3 | 3 | 3 | 104 | 3 | 8 | 20 |
| CC_3_2_3-pl_6 | 6 | 8 | 127 | 6 | 469 | 830 |
| CC_3_2_3-pl_7 | 8 | 42 | 325 | 7 | 3345 | 7256 |
| CC_3_3_3-pl_3 | 3 | 3 | 194 | 3 | 8 | 68 |
| CC_3_3_3-pl_5 | 10 | 745 | 14505 | 5 | 538 | 1347 |
| CC_3_3_3-pl_6 | 8 | 200 | 2970 | 6 | 1643 | 4670 |
| CC_3_3_3-pl_7 | 9 | 456 | 5932 | 7 | 12149 | 30566 |
| CoinBox-pl_2 | 2 | 2 | 84 | 2 | 2 | 8 |
| CoinBox-pl_5 | 5 | 75 | 279 | 5 | 102 | 118 |
| CoinBox-pl_7 | 7 | 1721 | 7102 | 7 | 2523 | 3214 |
| Grapevine_3-pl_2 | 2 | 4 | 114 | 2 | 6 | 22 |
| Grapevine_3-pl_5 | 5 | 823 | 2028 | |||
| Grapevine_3-pl_6 | 6 | 2118 | 4353 | |||
| Grapevine_3-pl_7 | 9 | 10414 | 144193 | 7 | 12008 | 19763 |
| Grapevine_4-pl_3 | 6 | 332 | 5342 | 3 | 40 | 612 |
| Grapevine_4-pl_4 | 7 | 190 | 10743 | 4 | 233 | 1429 |
| Grapevine_4-pl_5 | 6 | 90 | 6738 | 5 | 1449 | 7077 |
| Grapevine_4-pl_6 | 7 | 581 | 33904 | 6 | 4055 | 18234 |
| Grapevine_5-pl_2 | 3 | 3 | 398 | 2 | 8 | 815 |
| Grapevine_5-pl_3 | 6 | 17 | 3224 | 3 | 71 | 11056 |
| Grapevine_5-pl_5 | ||||||
| Grapevine_5-pl_6 | ||||||
| SC_R_10_10-pl_2 | 2 | 6 | 125 | 2 | 5 | 45 |
| SC_R_10_10-pl_3 | 3 | 10 | 109 | 3 | 17 | 88 |
| SC_R_10_10-pl_9 | 9 | 1248 | 7447 | 9 | 1112 | 6079 |
| SC_R_10_10-pl_13 | 13 | 17305 | 106761 | 13 | 22625 | 120270 |
| SC_R_10_10-pl_14 | 14 | 38714 | 239252 | 14 | 47255 | 239197 |
| SC_R_10_10-pl_17 | ||||||
| SC_4_1-pl_5 | 5 | 26 | 100 | 5 | 11 | 4 |
| SC_4_2-pl_5 | 5 | 549 | 431 | 5 | 22 | 24 |
| SC_4_2-pl_8 | 11 | 1116 | 942 | 8 | 318 | 258 |
| SC_4_3-pl_6 | 6 | 26 | 77 | 6 | 21 | 6 |
| SC_4_3-pl_8 | 8 | 75 | 112 | 8 | 66 | 16 |
| SC_4_4-pl_5 | 5 | 13 | 64 | 5 | 11 | 4 |
| SC_8_10-pl_8 | 12 | 802 | 3550 | 8 | 413 | 571 |
| SC_8_10-pl_9 | 12 | 1175 | 5247 | 9 | 1399 | 2267 |
| SC_8_10-pl_12 | 13 | 12336 | 42301 | 12 | 22486 | 34954 |
| SC_9_11-pl_4 | 4 | 4 | 65 | 4 | 6 | 8 |
| SC_9_11-pl_5 | 6 | 6 | 64 | 5 | 10 | 12 |
| SC_9_11-pl_7 | 7 | 9 | 75 | 7 | 32 | 30 |
| SC_9_11-pl_9 | 11 | 23 | 141 | 9 | 146 | 216 |
| SC_9_11-pl_10 | 12 | 30 | 199 | 10 | 355 | 590 |
| SC_9_11-pl_11 | 12 | 40 | 197 | 11 | 907 | 1617 |
| SC_10_8-pl_9 | ||||||
| SC_10_8-pl_14 | 16 | 19 | 83 | 14 | 51 | 40 |
| SC_10_10-pl_9 | 9 | 9 | 65 | 9 | 10 | 7 |
| SC_10_10-pl_10 | 11 | 11 | 72 | 10 | 12 | 9 |
| SC_10_10-pl_17 | 18 | 189 | 738 | 17 | 1075 | 1567 |
| () | 7 4 | 1559 5721 | 12150 38359 | 6 3 | 2462 7298 | 10425 34654 |
| - () | 7 4 | 64 296 | 1001 4635 | 6 3 | 287 1098 | 1068 4632 |
| () | 7 4 | 1559 5721 | 12150 38359 | 6 3 | 2495 7417 | 10671 35210 |
| - () | 7 4 | 64 296 | 1001 4635 | 6 4 | 242 1080 | 922 4870 |
| Solved Instances | 59/66 (89.39%) | 61/66 (92.42%) | ||||
G.2. Experimental Setup #2 – Knowledge Transfer
Experimental Setup #2 serves as the Knowledge Transfer experimental results.
To accomplish this, we trained the model - on the Collaboration and Communication and Grapevine domains, and tested it against all other domains. To evaluate this model against the other heuristics, we also included the Epistemic Gossip domain to expand the testing set for the various heuristics.
| Instance Name | GNN | BFS | ||||
| [ms] | [ms] | |||||
| CC_2_2_3-pl_4 | 4 | 14 | 129 | 4 | 17 | 35 |
| CC_2_2_3-pl_6 | 6 | 178 | 997 | 6 | 287 | 705 |
| CC_2_2_4-pl_5 | 5 | 94 | 1915 | 5 | 170 | 1262 |
| CC_2_3_4-pl_3 | 3 | 19 | 2510 | 3 | 6 | 446 |
| CC_2_3_4-pl_7 | 7 | 4153 | 341064 | 7 | 8907 | 171616 |
| CC_3_2_3-pl_4 | 4 | 47 | 418 | 4 | 30 | 91 |
| CC_3_2_3-pl_5 | 5 | 136 | 1331 | 5 | 178 | 549 |
| CC_3_3_3-pl_4 | 4 | 39 | 690 | 4 | 39 | 197 |
| Grapevine_3-pl_3 | 3 | 21 | 266 | 3 | 31 | 177 |
| Grapevine_3-pl_4 | 4 | 113 | 1863 | 4 | 160 | 1374 |
| Grapevine_4-pl_2 | 2 | 4 | 398 | 2 | 7 | 160 |
| Grapevine_5-pl_4 | 4 | 430 | 191371 | 4 | 616 | 120092 |
| () | 4 1 | 437 1126 | 45246 103440 | 4 1 | 871 2429 | 24725 55183 |
| - () | 4 1 | 75 126 | 1202 1651 | 4 1 | 101 178 | 556 1117 |
| () | 4 1 | 437 1126 | 45246 103440 | 4 1 | 871 2429 | 24725 55183 |
| - () | 4 1 | 75 126 | 1202 1651 | 4 1 | 101 178 | 556 1117 |
| Solved Instances | 12/12 (100.00%) | 12/12 (100.00%) | ||||
| Instance Name | GNN | BFS | ||||
| [ms] | [ms] | |||||
| CC_2_2_3-pl_3 | 3 | 9 | 123 | 3 | 9 | 12 |
| CC_2_2_3-pl_5 | 5 | 77 | 424 | 5 | 78 | 117 |
| CC_2_2_3-pl_7 | 7 | 558 | 2502 | 7 | 1065 | 1789 |
| CC_2_2_3-pl_8 | 8 | 2539 | 15968 | 8 | 2065 | 3288 |
| CC_2_2_4-pl_3 | 3 | 11 | 184 | 3 | 6 | 39 |
| CC_2_2_4-pl_4 | 4 | 23 | 332 | 4 | 24 | 155 |
| CC_2_2_4-pl_6 | 6 | 363 | 5495 | 6 | 822 | 2139 |
| CC_2_2_4-pl_7 | 7 | 2084 | 29932 | 7 | 4154 | 9578 |
| CC_2_3_4-pl_4 | 4 | 42 | 5574 | 4 | 29 | 1757 |
| CC_2_3_4-pl_5 | 5 | 127 | 19082 | 5 | 275 | 11988 |
| CC_2_3_4-pl_6 | 6 | 675 | 73622 | 6 | 1601 | 32067 |
| CC_3_2_3-pl_3 | 3 | 14 | 139 | 3 | 8 | 20 |
| CC_3_2_3-pl_6 | 6 | 367 | 3152 | 6 | 469 | 1287 |
| CC_3_2_3-pl_7 | 7 | 1548 | 14035 | 7 | 3345 | 10559 |
| CC_3_3_3-pl_3 | 3 | 15 | 368 | 3 | 8 | 77 |
| CC_3_3_3-pl_5 | 5 | 253 | 5069 | 5 | 538 | 1832 |
| CC_3_3_3-pl_6 | 6 | 1942 | 48781 | 6 | 1643 | 7964 |
| CC_3_3_3-pl_7 | 7 | 11726 | 424997 | 7 | 12149 | 46251 |
| CoinBox-pl_2 | 2 | 2 | 154 | 2 | 2 | 4 |
| CoinBox-pl_5 | 5 | 78 | 578 | 5 | 102 | 127 |
| CoinBox-pl_7 | 7 | 1793 | 13609 | 7 | 2523 | 4710 |
| Grapevine_3-pl_2 | 2 | 4 | 173 | 2 | 6 | 25 |
| Grapevine_3-pl_5 | 5 | 569 | 9163 | 5 | 823 | 3279 |
| Grapevine_3-pl_6 | 6 | 1621 | 23871 | 6 | 2118 | 6000 |
| Grapevine_3-pl_7 | 7 | 6781 | 132590 | 7 | 12008 | 26646 |
| Grapevine_4-pl_3 | 3 | 13 | 632 | 3 | 40 | 911 |
| Grapevine_4-pl_4 | 4 | 141 | 7522 | 4 | 233 | 2077 |
| Grapevine_4-pl_5 | 5 | 433 | 19923 | 5 | 1449 | 10105 |
| Grapevine_4-pl_6 | 6 | 6916 | 408771 | 6 | 4055 | 25563 |
| Grapevine_5-pl_2 | 2 | 4 | 1814 | 2 | 8 | 1184 |
| Grapevine_5-pl_3 | 3 | 27 | 13589 | 3 | 71 | 13375 |
| Grapevine_5-pl_5 | ||||||
| Grapevine_5-pl_6 | ||||||
| SC_R_10_10-pl_2 | 2 | 6 | 132 | 2 | 5 | 63 |
| SC_R_10_10-pl_3 | 3 | 10 | 123 | 3 | 17 | 142 |
| SC_R_10_10-pl_9 | 9 | 1248 | 7513 | 9 | 1112 | 9769 |
| SC_R_10_10-pl_13 | 13 | 17305 | 134216 | 13 | 22625 | 210204 |
| SC_R_10_10-pl_14 | 14 | 38714 | 278706 | 14 | 47255 | 483958 |
| SC_R_10_10-pl_17 | ||||||
| SC_4_1-pl_5 | 5 | 17 | 92 | 5 | 11 | 3 |
| SC_4_2-pl_5 | 5 | 30 | 141 | 5 | 22 | 20 |
| SC_4_2-pl_8 | 8 | 184 | 533 | 8 | 318 | 314 |
| SC_4_3-pl_6 | 6 | 21 | 96 | 6 | 21 | 6 |
| SC_4_3-pl_8 | 8 | 59 | 135 | 8 | 66 | 17 |
| SC_4_4-pl_5 | 5 | 17 | 71 | 5 | 11 | 4 |
| SC_8_10-pl_8 | 8 | 485 | 2414 | 8 | 413 | 635 |
| SC_8_10-pl_9 | 9 | 1772 | 10314 | 9 | 1399 | 2377 |
| SC_8_10-pl_12 | 12 | 35063 | 238864 | 12 | 22486 | 32874 |
| SC_9_11-pl_4 | 4 | 8 | 105 | 4 | 6 | 9 |
| SC_9_11-pl_5 | 5 | 11 | 110 | 5 | 10 | 12 |
| SC_9_11-pl_7 | 7 | 41 | 203 | 7 | 32 | 39 |
| SC_9_11-pl_9 | 9 | 237 | 1111 | 9 | 146 | 241 |
| SC_9_11-pl_10 | 10 | 657 | 3543 | 10 | 355 | 689 |
| SC_9_11-pl_11 | 11 | 1272 | 7150 | 11 | 907 | 1756 |
| SC_10_8-pl_9 | ||||||
| SC_10_8-pl_14 | 14 | 93 | 502 | 14 | 51 | 48 |
| SC_10_10-pl_9 | 9 | 10 | 143 | 9 | 10 | 9 |
| SC_10_10-pl_10 | 10 | 15 | 379 | 10 | 12 | 11 |
| SC_10_10-pl_17 | 17 | 942 | 7457 | 17 | 1075 | 1684 |
| () | 6 3 | 2527 7312 | 35931 91296 | 6 3 | 2729 7638 | 17633 69781 |
| - () | 6 4 | 288 1244 | 4116 13644 | 6 4 | 389 1410 | 1384 6943 |
| () | 6 3 | 2527 7312 | 35931 91296 | 6 3 | 2729 7638 | 17633 69781 |
| - () | 6 4 | 288 1244 | 4116 13644 | 6 4 | 389 1410 | 1384 6943 |
| Solved Instances | 55/59 (93.22%) | 55/59 (93.22%) | ||||
-EFP’s Individual Heuristics against Model -
| Instance Name | GNN | BFS | -EFP C_PG only | |||||||
| [ms] | [ms] | [ms] | Heur. | |||||||
| Assemble_B2-pl_5 | 5 | 14 | 140 | 5 | 14 | 47 | 5 | 5 | 97 | C_PG |
| Assemble_B4-pl_5 | 5 | 14 | 243 | 5 | 14 | 77 | 5 | 5 | 114 | C_PG |
| Assemble_B6-pl_5 | 5 | 14 | 760 | 5 | 14 | 461 | 5 | 5 | 1852 | C_PG |
| Assemble_B8-pl_5 | 5 | 14 | 31750 | 5 | 14 | 24112 | 5 | 5 | 147496 | C_PG |
| Assemble_B9-pl_5 | 5 | 14 | 585443 | 5 | 14 | 390542 | ||||
| Assemble_B10-pl_5 | ||||||||||
| Assemble_C-pl_5 | 5 | 14 | 132 | 5 | 14 | 49 | 5 | 5 | 87 | C_PG |
| CC_2_2_3-pl_3 | 3 | 9 | 62 | 3 | 9 | 17 | 5 | 5 | 64 | C_PG |
| CC_2_2_3-pl_5 | 5 | 76 | 362 | 5 | 77 | 176 | 5 | 5 | 49 | C_PG |
| CC_2_2_3-pl_7 | 7 | 451 | 2245 | 7 | 1052 | 2776 | ||||
| CC_2_2_3-pl_8 | 8 | 1934 | 11128 | 8 | 2045 | 5186 | ||||
| CC_2_2_4-pl_3 | 3 | 11 | 182 | 3 | 6 | 60 | 4 | 4 | 162 | C_PG |
| CC_2_2_4-pl_4 | 4 | 23 | 378 | 4 | 24 | 232 | ||||
| CC_2_2_4-pl_6 | 6 | 432 | 4769 | 6 | 824 | 5030 | ||||
| CC_2_2_4-pl_7 | 7 | 1911 | 21141 | 7 | 4171 | 16361 | ||||
| CC_2_3_4-pl_4 | 4 | 42 | 8451 | 4 | 29 | 2504 | ||||
| CC_2_3_4-pl_5 | 5 | 127 | 30013 | 5 | 277 | 18700 | ||||
| CC_2_3_4-pl_6 | 6 | 756 | 94259 | 6 | 1644 | 50540 | ||||
| CC_3_2_3-pl_3 | 3 | 14 | 123 | 3 | 8 | 49 | 5 | 5 | 143 | C_PG |
| CC_3_2_3-pl_6 | 6 | 368 | 4453 | 6 | 471 | 2205 | ||||
| CC_3_2_3-pl_7 | 7 | 1596 | 25601 | 7 | 3329 | 17005 | ||||
| CC_3_3_3-pl_3 | 3 | 15 | 699 | 3 | 8 | 180 | 5 | 5 | 590 | C_PG |
| CC_3_3_3-pl_5 | 5 | 253 | 6863 | 5 | 538 | 3136 | ||||
| CC_3_3_3-pl_6 | 6 | 1955 | 52929 | 6 | 1643 | 15139 | ||||
| CC_3_3_3-pl_7 | 7 | 11769 | 304660 | 7 | 12143 | 86997 | ||||
| CoinBox-pl_2 | 2 | 2 | 61 | 2 | 2 | 11 | 4 | 4 | 144 | C_PG |
| CoinBox-pl_5 | 5 | 77 | 425 | 5 | 101 | 377 | 5 | 5 | 139 | C_PG |
| CoinBox-pl_7 | 7 | 1816 | 18684 | 7 | 2490 | 7580 | ||||
| Grapevine_3-pl_2 | 2 | 4 | 111 | 2 | 6 | 60 | 3 | 3 | 100 | C_PG |
| Grapevine_3-pl_5 | 5 | 568 | 10866 | 5 | 821 | 4755 | ||||
| Grapevine_3-pl_6 | 6 | 1599 | 21958 | 6 | 2113 | 9928 | ||||
| Grapevine_3-pl_7 | 7 | 6561 | 92670 | 7 | 12014 | 61716 | ||||
| Grapevine_4-pl_3 | 3 | 13 | 824 | 3 | 40 | 1351 | ||||
| Grapevine_4-pl_4 | 4 | 141 | 7062 | 4 | 233 | 5094 | ||||
| Grapevine_4-pl_5 | 5 | 434 | 17301 | 5 | 1445 | 16673 | ||||
| Grapevine_4-pl_6 | 6 | 6790 | 310258 | 6 | 4066 | 40681 | ||||
| Grapevine_5-pl_2 | 2 | 4 | 1344 | 2 | 8 | 1906 | 3 | 3 | 1592 | C_PG |
| Grapevine_5-pl_3 | 3 | 27 | 17341 | 3 | 71 | 26331 | ||||
| Grapevine_5-pl_5 | ||||||||||
| Grapevine_5-pl_6 | ||||||||||
| SC_4_1-pl_5 | 5 | 17 | 73 | 5 | 11 | 21 | ||||
| SC_4_2-pl_5 | 5 | 15 | 215 | 5 | 21 | 49 | ||||
| SC_4_2-pl_8 | 8 | 339 | 2484 | 8 | 306 | 627 | ||||
| SC_4_3-pl_6 | 6 | 21 | 162 | 6 | 21 | 19 | ||||
| SC_4_3-pl_8 | 8 | 59 | 273 | 8 | 66 | 53 | ||||
| SC_4_4-pl_5 | 5 | 17 | 119 | 5 | 11 | 11 | ||||
| SC_8_10-pl_8 | 8 | 542 | 4565 | 8 | 388 | 1412 | ||||
| SC_8_10-pl_9 | 9 | 1660 | 20349 | 9 | 1332 | 4478 | ||||
| SC_8_10-pl_12 | 12 | 18541 | 187834 | 12 | 21604 | 35400 | ||||
| SC_9_11-pl_4 | 4 | 8 | 89 | 4 | 6 | 23 | 4 | 4 | 74 | C_PG |
| SC_9_11-pl_5 | 5 | 11 | 157 | 5 | 10 | 32 | 6 | 6 | 120 | C_PG |
| SC_9_11-pl_7 | 7 | 41 | 424 | 7 | 32 | 96 | 8 | 8 | 143 | C_PG |
| SC_9_11-pl_9 | 9 | 238 | 2497 | 9 | 146 | 439 | 12 | 12 | 285 | C_PG |
| SC_9_11-pl_10 | 10 | 663 | 6789 | 10 | 356 | 1364 | 13 | 13 | 272 | C_PG |
| SC_9_11-pl_11 | 11 | 1262 | 14034 | 11 | 913 | 2892 | 14 | 14 | 373 | C_PG |
| SC_10_8-pl_9 | ||||||||||
| SC_10_8-pl_14 | 14 | 93 | 544 | 14 | 51 | 127 | 15 | 24 | 160 | C_PG |
| SC_10_10-pl_9 | 9 | 10 | 151 | 9 | 10 | 35 | 9 | 9 | 92 | C_PG |
| SC_10_10-pl_10 | 10 | 15 | 223 | 10 | 12 | 29 | 10 | 10 | 87 | C_PG |
| SC_10_10-pl_17 | 17 | 946 | 6582 | 17 | 1078 | 3180 | ||||
| gossip_3_3_3 | 4 | 35 | 1257 | 4 | 41 | 170 | 5 | 5 | 179 | C_PG |
| gossip_3_3_6 | 2 | 3 | 195 | 2 | 4 | 50 | 4 | 4 | 152 | C_PG |
| gossip_3_3_9 | 6 | 356 | 3296 | 6 | 356 | 479 | 6 | 6 | 172 | C_PG |
| gossip_4_3_9 | 9 | 9 | 972 | C_PG | ||||||
| gossip_4_4_1 | 2 | 2 | 249 | 2 | 2 | 128 | 2 | 2 | 563 | C_PG |
| gossip_4_4_3 | 10 | 10 | 4486 | C_PG | ||||||
| gossip_4_4_7 | 6 | 6 | 1984 | C_PG | ||||||
| gossip_5_3_3 | 4 | 168 | 16236 | 4 | 179 | 2572 | 7 | 7 | 1400 | C_PG |
| gossip_5_3_5 | 0 | 0 | 114 | 0 | 0 | 6 | ||||
| gossip_5_3_7 | 8 | 8 | 1241 | C_PG | ||||||
| gossip_5_4_3 | 12 | 12 | 10860 | C_PG | ||||||
| gossip_5_4_4 | 1 | 1 | 267 | 1 | 1 | 68 | 1 | 1 | 207 | C_PG |
| gossip_5_4_9 | 16 | 16 | 13997 | C_PG | ||||||
| gossip_5_5_4 | 1 | 1 | 731 | 1 | 1 | 172 | 1 | 1 | 733 | C_PG |
| gossip_5_5_8 | 8 | 8 | 30331 | C_PG | ||||||
| () | 6 3 | 1014 2909 | 30556 92026 | 6 3 | 1230 3419 | 13625 50476 | 7 4 | 42 168 | 4928 30250 | |
| - () | 5 1 | 133 162 | 3791 4395 | 5 1 | 169 226 | 1518 1652 | 6 1 | 8 2 | 235 170 | |
| () | 6 3 | 1031 2930 | 31039 92684 | 6 3 | 1250 3443 | 13841 50851 | 7 4 | 46 179 | 5170 32344 | |
| - () | 5 1 | 133 162 | 3906 4419 | 5 1 | 180 230 | 1518 1652 | 6 1 | 8 3 | 164 107 | |
| Solved Instances | 64/75 () | 64/75 () | 37/75 () | |||||||
| Instance Name | GNN | BFS | -EFP L_PG only | |||||||
| [ms] | [ms] | [ms] | Heur. | |||||||
| Assemble_B2-pl_5 | 5 | 14 | 140 | 5 | 14 | 47 | 5 | 8 | 164 | L_PG |
| Assemble_B4-pl_5 | 5 | 14 | 243 | 5 | 14 | 77 | 5 | 8 | 218 | L_PG |
| Assemble_B6-pl_5 | 5 | 14 | 760 | 5 | 14 | 461 | 5 | 8 | 2995 | L_PG |
| Assemble_B8-pl_5 | 5 | 14 | 31750 | 5 | 14 | 24112 | 5 | 8 | 241692 | L_PG |
| Assemble_B9-pl_5 | 5 | 14 | 585443 | 5 | 14 | 390542 | ||||
| Assemble_B10-pl_5 | ||||||||||
| Assemble_C-pl_5 | 5 | 14 | 132 | 5 | 14 | 49 | 5 | 8 | 179 | L_PG |
| CC_2_2_3-pl_3 | 3 | 9 | 62 | 3 | 9 | 17 | 3 | 3 | 67 | L_PG |
| CC_2_2_3-pl_5 | 5 | 76 | 362 | 5 | 77 | 176 | 5 | 7 | 123 | L_PG |
| CC_2_2_3-pl_7 | 7 | 451 | 2245 | 7 | 1052 | 2776 | 11 | 35 | 676 | L_PG |
| CC_2_2_3-pl_8 | 8 | 1934 | 11128 | 8 | 2045 | 5186 | 9 | 16 | 443 | L_PG |
| CC_2_2_4-pl_3 | 3 | 11 | 182 | 3 | 6 | 60 | 3 | 3 | 180 | L_PG |
| CC_2_2_4-pl_4 | 4 | 23 | 378 | 4 | 24 | 232 | 4 | 4 | 196 | L_PG |
| CC_2_2_4-pl_6 | 6 | 432 | 4769 | 6 | 824 | 5030 | 7 | 19 | 1110 | L_PG |
| CC_2_2_4-pl_7 | 7 | 1911 | 21141 | 7 | 4171 | 16361 | 10 | 39 | 2253 | L_PG |
| CC_2_3_4-pl_4 | 4 | 42 | 8451 | 4 | 29 | 2504 | 4 | 4 | 1566 | L_PG |
| CC_2_3_4-pl_5 | 5 | 127 | 30013 | 5 | 277 | 18700 | 8 | 18 | 6684 | L_PG |
| CC_2_3_4-pl_6 | 6 | 756 | 94259 | 6 | 1644 | 50540 | 8 | 23 | 10986 | L_PG |
| CC_3_2_3-pl_3 | 3 | 14 | 123 | 3 | 8 | 49 | 3 | 3 | 126 | L_PG |
| CC_3_2_3-pl_6 | 6 | 368 | 4453 | 6 | 471 | 2205 | 6 | 13 | 493 | L_PG |
| CC_3_2_3-pl_7 | 7 | 1596 | 25601 | 7 | 3329 | 17005 | 12 | 57 | 2394 | L_PG |
| CC_3_3_3-pl_3 | 3 | 15 | 699 | 3 | 8 | 180 | 3 | 3 | 312 | L_PG |
| CC_3_3_3-pl_5 | 5 | 253 | 6863 | 5 | 538 | 3136 | 6 | 34 | 3583 | L_PG |
| CC_3_3_3-pl_6 | 6 | 1955 | 52929 | 6 | 1643 | 15139 | 9 | 28 | 3029 | L_PG |
| CC_3_3_3-pl_7 | 7 | 11769 | 304660 | 7 | 12143 | 86997 | 54 | 682 | 61837 | L_PG |
| CoinBox-pl_2 | 2 | 2 | 61 | 2 | 2 | 11 | 2 | 2 | 66 | L_PG |
| CoinBox-pl_5 | 5 | 77 | 425 | 5 | 101 | 377 | 5 | 10 | 763 | L_PG |
| CoinBox-pl_7 | 7 | 1816 | 18684 | 7 | 2490 | 7580 | 8 | 15 | 1319 | L_PG |
| Grapevine_3-pl_2 | 2 | 4 | 111 | 2 | 6 | 60 | 2 | 4 | 215 | L_PG |
| Grapevine_3-pl_5 | 5 | 568 | 10866 | 5 | 821 | 4755 | 53 | 493 | 34373 | L_PG |
| Grapevine_3-pl_6 | 6 | 1599 | 21958 | 6 | 2113 | 9928 | ||||
| Grapevine_3-pl_7 | 7 | 6561 | 92670 | 7 | 12014 | 61716 | ||||
| Grapevine_4-pl_3 | 3 | 13 | 824 | 3 | 40 | 1351 | 3 | 10 | 2114 | L_PG |
| Grapevine_4-pl_4 | 4 | 141 | 7062 | 4 | 233 | 5094 | 10 | 54 | 10507 | L_PG |
| Grapevine_4-pl_5 | 5 | 434 | 17301 | 5 | 1445 | 16673 | ||||
| Grapevine_4-pl_6 | 6 | 6790 | 310258 | 6 | 4066 | 40681 | ||||
| Grapevine_5-pl_2 | 2 | 4 | 1344 | 2 | 8 | 1906 | 2 | 4 | 4751 | L_PG |
| Grapevine_5-pl_3 | 3 | 27 | 17341 | 3 | 71 | 26331 | ||||
| Grapevine_5-pl_5 | ||||||||||
| Grapevine_5-pl_6 | ||||||||||
| SC_4_1-pl_5 | 5 | 17 | 73 | 5 | 11 | 21 | 5 | 6 | 45 | L_PG |
| SC_4_2-pl_5 | 5 | 15 | 215 | 5 | 21 | 49 | 5 | 6 | 83 | L_PG |
| SC_4_2-pl_8 | 8 | 339 | 2484 | 8 | 306 | 627 | ||||
| SC_4_3-pl_6 | 6 | 21 | 162 | 6 | 21 | 19 | ||||
| SC_4_3-pl_8 | 8 | 59 | 273 | 8 | 66 | 53 | ||||
| SC_4_4-pl_5 | 5 | 17 | 119 | 5 | 11 | 11 | 5 | 6 | 51 | L_PG |
| SC_8_10-pl_8 | 8 | 542 | 4565 | 8 | 388 | 1412 | ||||
| SC_8_10-pl_9 | 9 | 1660 | 20349 | 9 | 1332 | 4478 | ||||
| SC_8_10-pl_12 | 12 | 18541 | 187834 | 12 | 21604 | 35400 | ||||
| SC_9_11-pl_4 | 4 | 8 | 89 | 4 | 6 | 23 | 4 | 5 | 154 | L_PG |
| SC_9_11-pl_5 | 5 | 11 | 157 | 5 | 10 | 32 | 5 | 8 | 201 | L_PG |
| SC_9_11-pl_7 | 7 | 41 | 424 | 7 | 32 | 96 | 7 | 12 | 323 | L_PG |
| SC_9_11-pl_9 | 9 | 238 | 2497 | 9 | 146 | 439 | ||||
| SC_9_11-pl_10 | 10 | 663 | 6789 | 10 | 356 | 1364 | ||||
| SC_9_11-pl_11 | 11 | 1262 | 14034 | 11 | 913 | 2892 | ||||
| SC_10_8-pl_9 | ||||||||||
| SC_10_8-pl_14 | 14 | 93 | 544 | 14 | 51 | 127 | 14 | 14 | 289 | L_PG |
| SC_10_10-pl_9 | 9 | 10 | 151 | 9 | 10 | 35 | 9 | 9 | 170 | L_PG |
| SC_10_10-pl_10 | 10 | 15 | 223 | 10 | 12 | 29 | 10 | 10 | 212 | L_PG |
| SC_10_10-pl_17 | 17 | 946 | 6582 | 17 | 1078 | 3180 | ||||
| gossip_3_3_3 | 4 | 35 | 1257 | 4 | 41 | 170 | 4 | 10 | 380 | L_PG |
| gossip_3_3_6 | 2 | 3 | 195 | 2 | 4 | 50 | 2 | 3 | 152 | L_PG |
| gossip_3_3_9 | 6 | 356 | 3296 | 6 | 356 | 479 | 6 | 21 | 646 | L_PG |
| gossip_4_3_9 | 9 | 50 | 6289 | L_PG | ||||||
| gossip_4_4_1 | 2 | 2 | 249 | 2 | 2 | 128 | 2 | 2 | 470 | L_PG |
| gossip_4_4_3 | 8 | 45 | 26285 | L_PG | ||||||
| gossip_4_4_7 | 12 | 80 | 40528 | L_PG | ||||||
| gossip_5_3_3 | 4 | 168 | 16236 | 4 | 179 | 2572 | 6 | 27 | 7133 | L_PG |
| gossip_5_3_5 | 0 | 0 | 114 | 0 | 0 | 6 | 0 | 0 | 17 | L_PG |
| gossip_5_3_7 | 12 | 80 | 23547 | L_PG | ||||||
| gossip_5_4_3 | 10 | 69 | 38632 | L_PG | ||||||
| gossip_5_4_4 | 1 | 1 | 267 | 1 | 1 | 68 | 1 | 1 | 239 | L_PG |
| gossip_5_4_9 | 16 | 129 | 61902 | L_PG | ||||||
| gossip_5_5_4 | 1 | 1 | 731 | 1 | 1 | 172 | 1 | 1 | 689 | L_PG |
| gossip_5_5_8 | ||||||||||
| () | 6 3 | 1014 2909 | 30556 92026 | 6 3 | 1230 3419 | 13625 50476 | 8 10 | 41 111 | 11182 34781 | |
| - () | 5 1 | 133 162 | 3791 4395 | 5 1 | 169 226 | 1518 1652 | 6 5 | 12 24 | 1043 4275 | |
| () | 5 2 | 526 1732 | 13855 45504 | 5 2 | 684 1896 | 5755 14755 | 8 10 | 37 117 | 8472 35467 | |
| - () | 5 3 | 51 268 | 1584 7234 | 5 3 | 66 376 | 750 3492 | 5 5 | 9 14 | 645 2112 | |
| Solved Instances | 64/75 () | 64/75 () | 54/75 () | |||||||
| Instance Name | GNN | BFS | -EFP S_PG only | |||||||
| [ms] | [ms] | [ms] | Heur. | |||||||
| Assemble_B2-pl_5 | 5 | 14 | 140 | 5 | 14 | 47 | 5 | 8 | 176 | S_PG |
| Assemble_B4-pl_5 | 5 | 14 | 243 | 5 | 14 | 77 | 5 | 8 | 209 | S_PG |
| Assemble_B6-pl_5 | 5 | 14 | 760 | 5 | 14 | 461 | 5 | 8 | 3140 | S_PG |
| Assemble_B8-pl_5 | 5 | 14 | 31750 | 5 | 14 | 24112 | 5 | 8 | 239144 | S_PG |
| Assemble_B9-pl_5 | 5 | 14 | 585443 | 5 | 14 | 390542 | ||||
| Assemble_B10-pl_5 | ||||||||||
| Assemble_C-pl_5 | 5 | 14 | 132 | 5 | 14 | 49 | 5 | 8 | 178 | S_PG |
| CC_2_2_3-pl_3 | 3 | 9 | 62 | 3 | 9 | 17 | 3 | 3 | 89 | S_PG |
| CC_2_2_3-pl_5 | 5 | 76 | 362 | 5 | 77 | 176 | 5 | 5 | 92 | S_PG |
| CC_2_2_3-pl_7 | 7 | 451 | 2245 | 7 | 1052 | 2776 | 7 | 18 | 389 | S_PG |
| CC_2_2_3-pl_8 | 8 | 1934 | 11128 | 8 | 2045 | 5186 | ||||
| CC_2_2_4-pl_3 | 3 | 11 | 182 | 3 | 6 | 60 | 3 | 3 | 147 | S_PG |
| CC_2_2_4-pl_4 | 4 | 23 | 378 | 4 | 24 | 232 | 4 | 4 | 209 | S_PG |
| CC_2_2_4-pl_6 | 6 | 432 | 4769 | 6 | 824 | 5030 | 6 | 6 | 296 | S_PG |
| CC_2_2_4-pl_7 | 7 | 1911 | 21141 | 7 | 4171 | 16361 | 10 | 26 | 1542 | S_PG |
| CC_2_3_4-pl_4 | 4 | 42 | 8451 | 4 | 29 | 2504 | 4 | 4 | 1531 | S_PG |
| CC_2_3_4-pl_5 | 5 | 127 | 30013 | 5 | 277 | 18700 | 5 | 5 | 2390 | S_PG |
| CC_2_3_4-pl_6 | 6 | 756 | 94259 | 6 | 1644 | 50540 | 6 | 6 | 1731 | S_PG |
| CC_3_2_3-pl_3 | 3 | 14 | 123 | 3 | 8 | 49 | 3 | 3 | 167 | S_PG |
| CC_3_2_3-pl_6 | 6 | 368 | 4453 | 6 | 471 | 2205 | 6 | 6 | 284 | S_PG |
| CC_3_2_3-pl_7 | 7 | 1596 | 25601 | 7 | 3329 | 17005 | 10 | 27 | 1077 | S_PG |
| CC_3_3_3-pl_3 | 3 | 15 | 699 | 3 | 8 | 180 | 3 | 3 | 337 | S_PG |
| CC_3_3_3-pl_5 | 5 | 253 | 6863 | 5 | 538 | 3136 | 5 | 5 | 471 | S_PG |
| CC_3_3_3-pl_6 | 6 | 1955 | 52929 | 6 | 1643 | 15139 | 6 | 8 | 737 | S_PG |
| CC_3_3_3-pl_7 | 7 | 11769 | 304660 | 7 | 12143 | 86997 | ||||
| CoinBox-pl_2 | 2 | 2 | 61 | 2 | 2 | 11 | 2 | 2 | 72 | S_PG |
| CoinBox-pl_5 | 5 | 77 | 425 | 5 | 101 | 377 | 5 | 5 | 350 | S_PG |
| CoinBox-pl_7 | 7 | 1816 | 18684 | 7 | 2490 | 7580 | 8 | 9 | 705 | S_PG |
| Grapevine_3-pl_2 | 2 | 4 | 111 | 2 | 6 | 60 | 2 | 2 | 91 | S_PG |
| Grapevine_3-pl_5 | 5 | 568 | 10866 | 5 | 821 | 4755 | 7 | 15 | 867 | S_PG |
| Grapevine_3-pl_6 | 6 | 1599 | 21958 | 6 | 2113 | 9928 | 6 | 7 | 536 | S_PG |
| Grapevine_3-pl_7 | 7 | 6561 | 92670 | 7 | 12014 | 61716 | 11 | 26 | 2193 | S_PG |
| Grapevine_4-pl_3 | 3 | 13 | 824 | 3 | 40 | 1351 | 3 | 3 | 537 | S_PG |
| Grapevine_4-pl_4 | 4 | 141 | 7062 | 4 | 233 | 5094 | 4 | 4 | 553 | S_PG |
| Grapevine_4-pl_5 | 5 | 434 | 17301 | 5 | 1445 | 16673 | ||||
| Grapevine_4-pl_6 | 6 | 6790 | 310258 | 6 | 4066 | 40681 | ||||
| Grapevine_5-pl_2 | 2 | 4 | 1344 | 2 | 8 | 1906 | 2 | 2 | 1226 | S_PG |
| Grapevine_5-pl_3 | 3 | 27 | 17341 | 3 | 71 | 26331 | 3 | 3 | 1750 | S_PG |
| Grapevine_5-pl_5 | 7 | 13 | 10355 | S_PG | ||||||
| Grapevine_5-pl_6 | 6 | 7 | 12755 | S_PG | ||||||
| SC_4_1-pl_5 | 5 | 17 | 73 | 5 | 11 | 21 | 5 | 6 | 55 | S_PG |
| SC_4_2-pl_5 | 5 | 15 | 215 | 5 | 21 | 49 | 5 | 6 | 99 | S_PG |
| SC_4_2-pl_8 | 8 | 339 | 2484 | 8 | 306 | 627 | ||||
| SC_4_3-pl_6 | 6 | 21 | 162 | 6 | 21 | 19 | ||||
| SC_4_3-pl_8 | 8 | 59 | 273 | 8 | 66 | 53 | ||||
| SC_4_4-pl_5 | 5 | 17 | 119 | 5 | 11 | 11 | 5 | 6 | 68 | S_PG |
| SC_8_10-pl_8 | 8 | 542 | 4565 | 8 | 388 | 1412 | ||||
| SC_8_10-pl_9 | 9 | 1660 | 20349 | 9 | 1332 | 4478 | ||||
| SC_8_10-pl_12 | 12 | 18541 | 187834 | 12 | 21604 | 35400 | 13 | 13 | 996 | S_PG |
| SC_9_11-pl_4 | 4 | 8 | 89 | 4 | 6 | 23 | 4 | 4 | 127 | S_PG |
| SC_9_11-pl_5 | 5 | 11 | 157 | 5 | 10 | 32 | 6 | 7 | 220 | S_PG |
| SC_9_11-pl_7 | 7 | 41 | 424 | 7 | 32 | 96 | 7 | 11 | 337 | S_PG |
| SC_9_11-pl_9 | 9 | 238 | 2497 | 9 | 146 | 439 | 10 | 16 | 737 | S_PG |
| SC_9_11-pl_10 | 10 | 663 | 6789 | 10 | 356 | 1364 | 20 | 52 | 3455 | S_PG |
| SC_9_11-pl_11 | 11 | 1262 | 14034 | 11 | 913 | 2892 | 16 | 23 | 1294 | S_PG |
| SC_10_8-pl_9 | ||||||||||
| SC_10_8-pl_14 | 14 | 93 | 544 | 14 | 51 | 127 | 14 | 14 | 395 | S_PG |
| SC_10_10-pl_9 | 9 | 10 | 151 | 9 | 10 | 35 | 9 | 9 | 126 | S_PG |
| SC_10_10-pl_10 | 10 | 15 | 223 | 10 | 12 | 29 | 10 | 10 | 169 | S_PG |
| SC_10_10-pl_17 | 17 | 946 | 6582 | 17 | 1078 | 3180 | 22 | 22 | 763 | S_PG |
| gossip_3_3_3 | 4 | 35 | 1257 | 4 | 41 | 170 | 4 | 10 | 475 | S_PG |
| gossip_3_3_6 | 2 | 3 | 195 | 2 | 4 | 50 | 2 | 3 | 149 | S_PG |
| gossip_3_3_9 | 6 | 356 | 3296 | 6 | 356 | 479 | 6 | 21 | 696 | S_PG |
| gossip_4_3_9 | 9 | 50 | 6988 | S_PG | ||||||
| gossip_4_4_1 | 2 | 2 | 249 | 2 | 2 | 128 | 2 | 2 | 541 | S_PG |
| gossip_4_4_3 | 8 | 45 | 28848 | S_PG | ||||||
| gossip_4_4_7 | 12 | 80 | 42795 | S_PG | ||||||
| gossip_5_3_3 | 4 | 168 | 16236 | 4 | 179 | 2572 | 6 | 27 | 7701 | S_PG |
| gossip_5_3_5 | 0 | 0 | 114 | 0 | 0 | 6 | 0 | 0 | 18 | S_PG |
| gossip_5_3_7 | 12 | 80 | 23180 | S_PG | ||||||
| gossip_5_4_3 | 10 | 69 | 35226 | S_PG | ||||||
| gossip_5_4_4 | 1 | 1 | 267 | 1 | 1 | 68 | 1 | 1 | 198 | S_PG |
| gossip_5_4_9 | 16 | 129 | 58642 | S_PG | ||||||
| gossip_5_5_4 | 1 | 1 | 731 | 1 | 1 | 172 | 1 | 1 | 663 | S_PG |
| gossip_5_5_8 | ||||||||||
| () | 6 3 | 1014 2909 | 30556 92026 | 6 3 | 1230 3419 | 13625 50476 | 7 4 | 16 23 | 8085 31482 | |
| - () | 5 1 | 133 162 | 3791 4395 | 5 1 | 169 226 | 1518 1652 | 6 5 | 8 12 | 639 1501 | |
| () | 5 3 | 766 2638 | 12944 30900 | 5 3 | 1054 3336 | 6024 12442 | 6 4 | 10 9 | 5231 32153 | |
| - () | 5 4 | 86 403 | 2383 10062 | 5 4 | 97 511 | 1046 4309 | 5 4 | 7 8 | 456 793 | |
| Solved Instances | 64/75 () | 64/75 () | 62/75 () | |||||||
| Instance Name | GNN | BFS | -EFP SUB only | |||||||
| [ms] | [ms] | [ms] | Heur. | |||||||
| Assemble_B2-pl_5 | 5 | 14 | 140 | 5 | 14 | 47 | 5 | 10 | 274 | SUB |
| Assemble_B4-pl_5 | 5 | 14 | 243 | 5 | 14 | 77 | 5 | 10 | 287 | SUB |
| Assemble_B6-pl_5 | 5 | 14 | 760 | 5 | 14 | 461 | 5 | 10 | 921 | SUB |
| Assemble_B8-pl_5 | 5 | 14 | 31750 | 5 | 14 | 24112 | 5 | 10 | 40687 | SUB |
| Assemble_B9-pl_5 | 5 | 14 | 585443 | 5 | 14 | 390542 | 5 | 10 | 594127 | SUB |
| Assemble_B10-pl_5 | ||||||||||
| Assemble_C-pl_5 | 5 | 14 | 132 | 5 | 14 | 49 | 5 | 10 | 272 | SUB |
| CC_2_2_3-pl_3 | 3 | 9 | 62 | 3 | 9 | 17 | 4 | 4 | 77 | SUB |
| CC_2_2_3-pl_5 | 5 | 76 | 362 | 5 | 77 | 176 | 5 | 6 | 84 | SUB |
| CC_2_2_3-pl_7 | 7 | 451 | 2245 | 7 | 1052 | 2776 | 11 | 42 | 416 | SUB |
| CC_2_2_3-pl_8 | 8 | 1934 | 11128 | 8 | 2045 | 5186 | ||||
| CC_2_2_4-pl_3 | 3 | 11 | 182 | 3 | 6 | 60 | 3 | 3 | 188 | SUB |
| CC_2_2_4-pl_4 | 4 | 23 | 378 | 4 | 24 | 232 | 5 | 10 | 404 | SUB |
| CC_2_2_4-pl_6 | 6 | 432 | 4769 | 6 | 824 | 5030 | 9 | 18 | 634 | SUB |
| CC_2_2_4-pl_7 | 7 | 1911 | 21141 | 7 | 4171 | 16361 | 7 | 18 | 1263 | SUB |
| CC_2_3_4-pl_4 | 4 | 42 | 8451 | 4 | 29 | 2504 | 5 | 10 | 4796 | SUB |
| CC_2_3_4-pl_5 | 5 | 127 | 30013 | 5 | 277 | 18700 | 5 | 8 | 7774 | SUB |
| CC_2_3_4-pl_6 | 6 | 756 | 94259 | 6 | 1644 | 50540 | 9 | 22 | 8670 | SUB |
| CC_3_2_3-pl_3 | 3 | 14 | 123 | 3 | 8 | 49 | 3 | 3 | 137 | SUB |
| CC_3_2_3-pl_6 | 6 | 368 | 4453 | 6 | 471 | 2205 | 7 | 8 | 188 | SUB |
| CC_3_2_3-pl_7 | 7 | 1596 | 25601 | 7 | 3329 | 17005 | 7 | 13 | 578 | SUB |
| CC_3_3_3-pl_3 | 3 | 15 | 699 | 3 | 8 | 180 | 3 | 3 | 471 | SUB |
| CC_3_3_3-pl_5 | 5 | 253 | 6863 | 5 | 538 | 3136 | 6 | 8 | 789 | SUB |
| CC_3_3_3-pl_6 | 6 | 1955 | 52929 | 6 | 1643 | 15139 | 6 | 8 | 770 | SUB |
| CC_3_3_3-pl_7 | 7 | 11769 | 304660 | 7 | 12143 | 86997 | ||||
| CoinBox-pl_2 | 2 | 2 | 61 | 2 | 2 | 11 | 2 | 2 | 89 | SUB |
| CoinBox-pl_5 | 5 | 77 | 425 | 5 | 101 | 377 | 7 | 9 | 223 | SUB |
| CoinBox-pl_7 | 7 | 1816 | 18684 | 7 | 2490 | 7580 | ||||
| Grapevine_3-pl_2 | 2 | 4 | 111 | 2 | 6 | 60 | 2 | 2 | 94 | SUB |
| Grapevine_3-pl_5 | 5 | 568 | 10866 | 5 | 821 | 4755 | 5 | 6 | 322 | SUB |
| Grapevine_3-pl_6 | 6 | 1599 | 21958 | 6 | 2113 | 9928 | 6 | 7 | 516 | SUB |
| Grapevine_3-pl_7 | 7 | 6561 | 92670 | 7 | 12014 | 61716 | ||||
| Grapevine_4-pl_3 | 3 | 13 | 824 | 3 | 40 | 1351 | 3 | 3 | 475 | SUB |
| Grapevine_4-pl_4 | 4 | 141 | 7062 | 4 | 233 | 5094 | 4 | 4 | 426 | SUB |
| Grapevine_4-pl_5 | 5 | 434 | 17301 | 5 | 1445 | 16673 | 6 | 9 | 1246 | SUB |
| Grapevine_4-pl_6 | 6 | 6790 | 310258 | 6 | 4066 | 40681 | 6 | 7 | 2320 | SUB |
| Grapevine_5-pl_2 | 2 | 4 | 1344 | 2 | 8 | 1906 | 2 | 2 | 1491 | SUB |
| Grapevine_5-pl_3 | 3 | 27 | 17341 | 3 | 71 | 26331 | 3 | 3 | 1811 | SUB |
| Grapevine_5-pl_5 | 5 | 6 | 4094 | SUB | ||||||
| Grapevine_5-pl_6 | 7 | 11 | 15754 | SUB | ||||||
| SC_4_1-pl_5 | 5 | 17 | 73 | 5 | 11 | 21 | ||||
| SC_4_2-pl_5 | 5 | 15 | 215 | 5 | 21 | 49 | ||||
| SC_4_2-pl_8 | 8 | 339 | 2484 | 8 | 306 | 627 | ||||
| SC_4_3-pl_6 | 6 | 21 | 162 | 6 | 21 | 19 | ||||
| SC_4_3-pl_8 | 8 | 59 | 273 | 8 | 66 | 53 | ||||
| SC_4_4-pl_5 | 5 | 17 | 119 | 5 | 11 | 11 | ||||
| SC_8_10-pl_8 | 8 | 542 | 4565 | 8 | 388 | 1412 | ||||
| SC_8_10-pl_9 | 9 | 1660 | 20349 | 9 | 1332 | 4478 | ||||
| SC_8_10-pl_12 | 12 | 18541 | 187834 | 12 | 21604 | 35400 | ||||
| SC_9_11-pl_4 | 4 | 8 | 89 | 4 | 6 | 23 | 4 | 4 | 59 | SUB |
| SC_9_11-pl_5 | 5 | 11 | 157 | 5 | 10 | 32 | 6 | 6 | 66 | SUB |
| SC_9_11-pl_7 | 7 | 41 | 424 | 7 | 32 | 96 | 8 | 8 | 96 | SUB |
| SC_9_11-pl_9 | 9 | 238 | 2497 | 9 | 146 | 439 | 12 | 12 | 176 | SUB |
| SC_9_11-pl_10 | 10 | 663 | 6789 | 10 | 356 | 1364 | 13 | 13 | 206 | SUB |
| SC_9_11-pl_11 | 11 | 1262 | 14034 | 11 | 913 | 2892 | 14 | 14 | 239 | SUB |
| SC_10_8-pl_9 | ||||||||||
| SC_10_8-pl_14 | 14 | 93 | 544 | 14 | 51 | 127 | 16 | 33 | 205 | SUB |
| SC_10_10-pl_9 | 9 | 10 | 151 | 9 | 10 | 35 | 9 | 10 | 51 | SUB |
| SC_10_10-pl_10 | 10 | 15 | 223 | 10 | 12 | 29 | 11 | 12 | 77 | SUB |
| SC_10_10-pl_17 | 17 | 946 | 6582 | 17 | 1078 | 3180 | ||||
| gossip_3_3_3 | 4 | 35 | 1257 | 4 | 41 | 170 | 4 | 10 | 250 | SUB |
| gossip_3_3_6 | 2 | 3 | 195 | 2 | 4 | 50 | 2 | 3 | 100 | SUB |
| gossip_3_3_9 | 6 | 356 | 3296 | 6 | 356 | 479 | 6 | 21 | 408 | SUB |
| gossip_4_3_9 | 9 | 50 | 4497 | SUB | ||||||
| gossip_4_4_1 | 2 | 2 | 249 | 2 | 2 | 128 | 2 | 2 | 343 | SUB |
| gossip_4_4_3 | 8 | 45 | 20623 | SUB | ||||||
| gossip_4_4_7 | 12 | 80 | 54239 | SUB | ||||||
| gossip_5_3_3 | 4 | 168 | 16236 | 4 | 179 | 2572 | 6 | 27 | 6392 | SUB |
| gossip_5_3_5 | 0 | 0 | 114 | 0 | 0 | 6 | 0 | 0 | 15 | SUB |
| gossip_5_3_7 | 12 | 80 | 18720 | SUB | ||||||
| gossip_5_4_3 | 10 | 69 | 23396 | SUB | ||||||
| gossip_5_4_4 | 1 | 1 | 267 | 1 | 1 | 68 | 1 | 1 | 196 | SUB |
| gossip_5_4_9 | 16 | 129 | 33129 | SUB | ||||||
| gossip_5_5_4 | 1 | 1 | 731 | 1 | 1 | 172 | 1 | 1 | 596 | SUB |
| gossip_5_5_8 | ||||||||||
| () | 6 3 | 1014 2909 | 30556 92026 | 6 3 | 1230 3419 | 13625 50476 | 6 4 | 16 23 | 14772 77443 | |
| - () | 5 1 | 133 162 | 3791 4395 | 5 1 | 169 226 | 1518 1652 | 6 4 | 9 9 | 573 2003 | |
| () | 5 3 | 414 1040 | 26116 91820 | 5 3 | 504 985 | 13305 54917 | 6 3 | 10 8 | 13646 83133 | |
| - () | 5 3 | 77 354 | 2742 13014 | 5 3 | 89 433 | 1193 4899 | 5 4 | 8 7 | 370 605 | |
| Solved Instances | 64/75 () | 64/75 () | 58/75 () | |||||||
Appendix H Overhead of Single-Query GNN-Based Estimation
Although our current implementation evaluates the GNN heuristic one state at a time, we can estimate the potential impact of batching using aggregate runtime statistics.
From Table 4, BFS requires ms (time per expanded node), whereas the GNN-based configuration requires ms per node. The difference, ms, represents the effective overhead of a single-query GNN evaluation.
Assuming that batching affects only this scoring component, we model the per-node cost under batch size as
where ms denotes the planner-dominated per-node cost. For , this yields ms, matching the observed single-query runtime.
The total projected runtime under batching is then
where is the number of nodes expanded by GNN.
Figure 2 reports the resulting estimates for . The results show that batching rapidly amortizes the GNN overhead: with , the projected runtime decreases from ms to approximately ms, approaching the planner-dominated regime. This supports our choice of expanded nodes as the primary evaluation metric, as heuristic informativeness remains invariant under such engineering optimizations.
Finally, we quantify the relative speed-up induced by batching. Given the projected runtime , we define the speed-up with respect to single-query inference () as
Note that , with diminishing returns as approaches the planner-dominated baseline .